### v1

# Unlearning Isn’t Deletion: Investigating Reversibility of Machine Unlearning in LLMs

Abstract
Unlearning in large language models (LLMs) is intended to remove the influence of specific data, yet current evaluations rely heavily on token-level metrics such as accuracy and perplexity. We show that these metrics can be misleading: models often appear to forget, but their original behavior can be rapidly restored with minimal fine-tuning, revealing that unlearning may obscure information rather than erase it. To diagnose this phenomenon, we introduce a representation-level evaluation framework using PCA-based similarity and shift, centered kernel alignment, and Fisher information. Applying this toolkit across six unlearning methods, three domains (text, code, math), and two open-source LLMs, we uncover a critical distinction between reversible and irreversible forgetting. In reversible cases, models suffer token-level collapse yet retain latent features; in irreversible cases, deeper representational damage occurs. We further provide a theoretical account linking shallow weight perturbations near output layers to misleading unlearning signals, and show that reversibility is modulated by task type and hyperparameters. Our findings reveal a fundamental gap in current evaluation practices and establish a new diagnostic foundation for trustworthy unlearning in LLMs. We provide a unified toolkit for analyzing LLM representation changes under unlearning and relearning: https://github.com/XiaoyuXU1/Representational_Analysis_Tools.git.
1 Introduction
Large language models (LLMs), trained on massive corpora, have achieved remarkable success across diverse tasks, yet their capacity to memorize training snippets poses acute ethical, legal, and security risks. Memorization can unintentionally disclose sensitive, harmful, or copyrighted text [25; 13; 30], conflicting with emerging regulations, such as the EU’s Right to be Forgotten [9].
Machine unlearning seeks to mitigate this threat by making a model act as though specified data were never seen [2]. Numerous methods have been proposed for LLMs [33; 12; 6; 26; 18; 17; 19; 20; 31], their success usually judged by token-level metrics, such as accuracy or perplexity.
However, a pivotal question remains largely unexplored: Does LLM unlearning truly erase information, or merely suppress it, poised to “resurface” at the slightest nudge?
Empirically, many unlearning methods may just appear effective: after unlearning, a model shows near-zero accuracy or high perplexity on the forget set, but a brief fine-tuning step (even on unrelated data) can quickly restore its original behaviour (see Figure 1). This exposes a huge gap between surface-level metrics and the model’s internal state [21; 23], casting doubt on compliance and safety claims. If information is recoverable via simple fine-tuning (or relearning), can we truly claim that it has been “forgotten?” What looks like memory loss may, in fact, be a shallow perturbation.
In this work, we perform the first systematic analysis of reversibility of LLM unlearning, covering both single-shot and continual settings. The continual setting allows multiple unlearning requests over time, which we believe is a more prominent and realistic setting where the deployment of such unlearning strategies are often facing a dynamic environment. We show that standard token-level metrics prove insufficient—they can collapse even when the underlying representations remain intact. To probe deeper, we introduce a diagnostic toolkit for representational analysis, featuring PCA subspace similarity and shift [36], centered kernel alignment (CKA) [15], and Fisher information [3]. Our toolkit uncovers two distinct regimes of unlearning:
1) Reversible (catastrophic) forgetting: performance collapses, but feature subspaces are largely preserved, enabling rapid recovery, and 2) Irreversible (catastrophic) forgetting: collapse coincides with substantial representational drift, making recovery difficult/impossible. Surprisingly, both yield similar results under token-level metrics, underscoring the need for deeper representational analysis.
We further show that the transition between reversible and irreversible forgetting depends not only on the volume of unlearning requests but also on hyperparameters such as learning rate. Modest weight perturbations—especially near the output layer—can lead to token-level distortions without altering feature geometry, making “forgotten” knowledge easily recoverable.
Hence, evaluating the effectiveness of LLM unlearning must go beyond superficial token-level metrics (e.g., forget accuracy declines). In safety- and privacy-critical settings, unlearning should be judged by its ability to achieve genuine erasure rather than simply representational collapse.
Figure 1: (a) Token-level accuracy and CKA subspaces of reversible (top) vs. irreversible (bottom) catastrophic forgetting due to continual unlearning then relearning, (b) Our four diagnostic tools
Contributions. We summarize our main contributions as follows:
  • We present the first systematic study of reversibility in both single and continual LLM unlearning, using a feature-space toolkit, including PCA similarity, PCA shift, CKA, and Fisher information. Our analysis distinguishes reversible from irreversible (catastrophic) forgetting.
  • We conduct extensive experiments with six unlearning methods (GA [33], NPO [35], and RLabel, with their variants) across three datasets (arXiv papers, GitHub code [33], and NuminaMath-1.5 [16]) on Yi-6B [34] and Qwen-2.5-7B [32]. Our results show that standard token-level metrics (e.g., accuracy, perplexity, MIA susceptibility [27]) fail to capture true forgetting behavior.
  • We theoretically analyze weight perturbations to explain how widespread vs. localized parameter changes relate to (ir)reversible forgetting. Small perturbations near the logits can distort token-level metrics despite intact features, hence leading to misleading assessments.
  • Based on our findings and extra preliminary results, we propose several future directions, including using unlearning as a complementary form of data augmentation, and designing more robust unlearning algorithms that achieve genuine forgetting while avoiding representational collapse.
2 Preliminaries
LLM unlearning seeks to enhance privacy, improve safety, and reduce bias [33; 12; 26; 18; 17; 19]. Most work adopts the single-unlearning paradigm: given a training corpus $\mathcal{D}$ and a designated forget set $\mathcal{D}_{f}\subseteq\mathcal{D}$, a model $\mathcal{M}$ is first trained on $\mathcal{D}$ with algorithm $\mathcal{A}$. An unlearning procedure $\mathcal{U}$ then transforms $\mathcal{M}$ into an unlearned model $\mathcal{M}_{f}$ that should behave as if it had never encountered $\mathcal{D}_{f}$.
Ideally, $\mathcal{U}$ should produce a model statistically indistinguishable from one retrained on the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$: $\mathcal{M}_{f}=\mathcal{U}\!\bigl{(}\mathcal{M},\mathcal{D}_{f}\bigr{)}\approx\mathcal{M}_{r}=\mathcal{A}\!\bigl{(}\mathcal{M},\mathcal{D}_{r}\bigr{)}$. While current methods strike a good forget-utility balance in controlled settings [1; 8] (e.g., a fixed $\mathcal{D}_{f}$ or just a single removal request), they seldom address the practical need for continual unlearning, i.e., data owners can submit removal requests sequentially over time [1]. Let the successive forget sets be $\mathcal{D}_{f}^{(1)},\mathcal{D}_{f}^{(2)},\ldots,\mathcal{D}_{f}^{(t)}$ (whose union is $\mathcal{D}_{f}$); the retain set after $t$ rounds is $\mathcal{D}_{r}^{(t)}$. The model is then updated recursively: $\mathcal{M}_{f}^{(t+1)}=\mathcal{U}(\mathcal{M}_{f}^{(t)},\mathcal{D}_{f}^{(t+1)})$, which should be similar to $\mathcal{M}_{r}=\mathcal{A}\!\bigl{(}\mathcal{M},\mathcal{D}_{r}^{(t+1)}\bigr{)}$ at any time $t$.
Retraining LLMs is prohibitively costly, so most studies rely on empirical proxies rather than formal statistically-indistinguishable guarantees [6; 24; 18; 7]. Evaluations track forget quality on the forget set and utility on the retain set, aiming to preserve both [24]. While single unlearning often yields modest declines, it is fragile: brief fine-tuning—even on benign, unrelated data—swiftly revives the “forgotten” knowledge [1; 23; 21]. The issue worsens under continual unlearning, where each round begins from an already degraded model, ultimately triggering catastrophic forgetting—a wholesale performance collapse [1; 28]. Prior work notes this risk but does not examine its root causes.
We hypothesize that collapse does not imply true erasure; the knowledge may remain latent in the feature space. This insight leads us to distinguish two regimes of (catastrophic) forgetting:
Definition 1 (Reversible (Catastrophic) Forgetting).
Let $\theta_{0}$ denote the initial model parameters, $\mathcal{D}_{f}$ the forget set, and $\mathcal{T}$ an evaluation task with metric $E(\cdot,\mathcal{T})$. Unlearning $\mathcal{D}_{f}$ transforms the model to $\theta_{u}$. If subsequent relearning on $\mathcal{D}_{f}$ (or an equivalent reconstruction set) produces parameters $\theta_{r}$ s.t.
$E(\theta_{u},\mathcal{T})\;\ll\;E(\theta_{r},\mathcal{T})\;\approx\;E(\theta_{0},\mathcal{T}),$
the temporary performance collapse is fully reversible; we call it reversible catastrophic forgetting. When the initial degradation is modest (e.g., single unlearning), we simply call it reversible forgetting.
Definition 2 (Irreversible (Catastrophic) Forgetting).
Using the notations of Definition 1, if
$E(\theta_{u},\mathcal{T})\;\approx\;E(\theta_{r},\mathcal{T})\;\ll\;E(\theta_{0},\mathcal{T}),$
we observe irreversible catastrophic forgetting: the collapse (i.e., weight perturbation) is irreversible. We refine this to irreversible forgetting, which further requires that the irreversible degradation be restricted to the forget set; performance on the retain set and on unrelated data must remain near their original levels. This condition distinguishes targeted erasure from global model failure or collapse.
To distinguish our setting from a full retrain model, where reversibility is trivially achievable, we introduce the following restriction on the relearning phase.
Relearning Restriction. After unlearning, we briefly fine-tune $\theta_{u}$ on a small relearning set—either the cumulative forget set $\mathcal{D}_{f}=\bigcup_{t}\mathcal{D}_{f}^{(t)}$, its a distribution-similar retain set $\mathcal{D}_{r}^{(t)}$, or an unrelated out-of-distribution corpus—to obtain $\theta_{r}$ without ever revisiting the full pre-training data.
3 Token-Level Evaluation
3.1 Experiment setup
Models and Datasets. We conduct experiments on two open-source models, Yi-6B [34] and Qwen-2.5-7B [32]. To ensure the generality of our findings, we use two dataset types: i) simple tasks—arXiv abstracts and GitHub code from [33], and ii) a complex task—NuminaMath-1.5, a recent benchmark for mathematical reasoning [16]. All experiments were run on NVIDIA H100 GPUs.
Unlearning algorithms We compare six canonical methods grouped into three families.
1) Gradient-Ascent (GA) family. The unified goal is $\mathcal{L}\;=\;\mathcal{L}_{\text{forget}}\!\bigl{(}\mathcal{D}_{f}\bigr{)}+\;\lambda\,\mathcal{L}_{\text{retain}}\!\bigl{(}\mathcal{D}_{r}\bigr{)},$ where $\mathcal{L}_{\mathrm{forget}}$ maximizes the loss on the forget set via GA, $\mathcal{L}_{\mathrm{retain}}$ (optional) preserves utility on the retain set, and $\lambda>0$ balances the two. Choices for $\mathcal{L}_{\mathrm{retain}}$ give three variants: i) GA ($\mathcal{L}_{\mathrm{retain}}=0$), ii) GA+GD (standard cross-entropy on $\mathcal{D}{r}$), and iii) GA+KL (KL divergence to the reference model on $\mathcal{D}{r}$[33].
2) Negative Preference Optimization (NPO) family. GA is replaced by an NPO loss that penalizes agreement with the forget set [35]: $\mathcal{L}\;=\;\mathcal{L}_{\text{NPO}}\!\bigl{(}\mathcal{D}_{f}\bigr{)}+\;\lambda\,\mathcal{L}_{\text{retain}}\!\bigl{(}\mathcal{D}_{r}\bigr{)},$ Variants mirror those above: NPO ($\mathcal{L}_{\mathrm{retain}}=0$) and NPO+KL (retain-set KL regularization).
3) Random Label (RLabel). To mimic a model that never saw $\mathcal{D}_{f}$, true labels are replaced with random ones: $\mathcal{L}\;=\;\mathcal{L}_{\text{RLabel}}\!\bigl{(}\mathcal{D}_{f}\bigr{)},$ inducing near-uniform predictions without GA/negative rewards [33].
Unlearning Scenario We consider two standard settings: i) Single unlearning: A trained model $\mathcal{M}$ receives exactly one request to remove $\mathcal{D}_{f}\subset\mathcal{D}$, producing $\mathcal{M}_{f}=\mathcal{U}(\mathcal{M},\mathcal{D}_{f})$, and ii) Continual unlearning: The model processes a stream of requests ${\mathcal{D}_{f}^{(1)},\dots,\mathcal{D}_{f}^{(k)}}$, updated iteratively by $\mathcal{M}^{(i+1)}=\mathcal{U}(\mathcal{M}^{(t)},\mathcal{D}_{f}^{(i+1)})$ with $\mathcal{M}^{(0)}=\mathcal{M}$ and $\bigcup_{i=1}^{n}\mathcal{D}_{f}^{(t)}=\mathcal{D}_{f}$. This mirrors real-world, incremental removal demands while maintaining parity with the single-step budget.
For the simple tasks, we benchmark all six algorithms: GA, GA + GD, GA + KL, RLabel, NPO, and NPO+KL. For the complex one, lacking a clearly defined retain set, we use GA, NPO, and RLabel.
Evaluation Metrics For single unlearning (simple tasks only), we report: forget-set accuracy (F.Acc), retain-set accuracy (R.Acc), and privacy leakage via min-$k\%$-prob MIA AUC [27].
[…]
Relearning setting. To gauge how readily forgotten knowledge re-emerges, each unlearning run is followed by a controlled relearning phase. Single unlearning: We fine-tune on the whole forget set $\mathcal{D}_{f}$ once, producing a single-step relearned model. Continual unlearning: For settings that trigger catastrophic collapse, we fine-tune with three cases: i) the cumulative forget set $\bigcup_{i}\mathcal{D}_{f}^{(t)}$, ii) the corresponding retain set $\mathcal{D}_{r}^{(t)}$, and iii) an unrelated auxiliary corpus. These progressively relax assumptions about access to the forgotten content, revealing the recovery potential in each case.
Hyperparameter Configuration. To comprehensively evaluate the effects of unlearning, we design multiple hyperparameter configurations that vary both the learning rate and the number of unlearning requests. For single unlearning we sweep the learning rate over $LR\in\{3,4,5\}\!\times\!10^{-6}$ while fixing the request count to $N=1$. For continual unlearning we vary both knobs: on the simple task (Yi-6B) we test $LR\in\{3,5\}\!\times\!10^{-6}\cup\{3\times 10^{-5}\}$ with $N\in\{6\to 100\}$; on the complex task (Qwen-2.5-7B) we use $LR\in\{3,5\}\!\times\!10^{-6}$ and $3\times 10^{-5}$ together with $N\in\{6\to 100\}$. All runs adopt the optimizer settings of [1]: AdamW [22] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\varepsilon=10^{-8}$), a cosine schedule with 10% warm-up followed by decay to 10% of peak, weight decay 0.1, and gradient clipping at 1.0.
Table 1: Yi-6B: MIA / F.Acc / R.Acc (%) simple task using three LRs under single unlearning
Phase Method LR=3$\times$10-6 LR=4$\times$10-6 LR=5$\times$10-6
MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc
Original 70.9 78.9 65.5 70.9 78.9 65.5 70.9 78.9 65.5
Unlearn GA 45.5 65.4 54.0 43.8 62.4 52.3 41.2 60.3 50.9
GA+GD 65.4 75.1 64.6 58.2 73.8 65.8 55.3 68.5 63.5
GA+KL 48.9 71.0 58.5 47.6 70.6 58.1 44.8 68.4 55.4
NPO 67.2 76.2 64.7 65.2 75.8 62.8 62.2 75.2 62.7
NPO+KL 66.5 76.3 64.8 67.2 76.4 63.2 64.5 75.6 61.2
RLabel 69.6 77.7 64.7 69.2 76.5 64.5 68.7 75.4 63.3
Relearn GA 67.2 76.6 65.2 68.6 77.6 62.8 67.6 76.9 65.5
GA+GD 68.6 77.0 65.3 68.8 76.9 65.3 68.8 77.2 65.3
GA+KL 67.9 77.6 65.3 68.3 75.5 65.2 67.7 77.2 65.2
NPO 68.2 77.1 65.3 68.2 77.2 65.2 68.3 77.0 65.1
NPO+KL 68.9 77.1 65.3 67.9 76.3 63.0 68.6 76.9 65.2
RLabel 68.3 78.8 65.6 68.9 76.4 65.3 68.8 78.9 65.2
3.2 Token-Level Evaluation Results
We report quantitative results on both single unlearning and continual unlearning settings using Yi-6B and Qwen-2.5-7B across multiple configurations (Tables 13). For completeness, detailed experimental results and additional settings are provided in Appendix A.4.
Single Unlearning. For Yi-6B under single unlearning, Table 1 shows that all six unlearning methods reduce MIA and F.Acc, indicating some level of unlearning. The changes on retain set are modest: R.Acc only drops 2–5% for most methods, and MIA decreases by less than 30 points in most cases. Importantly, the relearned models often recover original performance—both GA+KL and RLabel restore R.Acc to near 65.0%, and F.Acc rebounds above 77%. These results suggest that under single unlearning, most methods achieve seemingly successful forgetting at the token level, but as we show in Section 4.2 to Section 4.2, the underlying representation changes are minimal—indicating the phenomenon of reversible forgetting.
Continual Unlearning. By examining post-relearning recoverability in Table 2 and Table 3, we identify two distinct forms of catastrophic forgetting. When the model regains both utility (e.g., F.Acc, R.Acc) and privacy (e.g., MIA AUC) to levels near or exceeding the original after relearning, we classify the behavior as reversible catastrophic forgetting. This suggests that the underlying representational structure remains intact, enabling efficient recovery via lightweight retraining. Such reversibility is consistently observed in methods like NPO and NPO+KL, particularly under low learning rates or small removal batches.
Conversely, when relearning fails to restore utility—reflected in persistently low F.Acc and R.Acc despite partial MIA recovery—we categorize it as irreversible catastrophic forgetting. This scenario frequently arises with methods like GA and RLabel under aggressive hyperparameters (e.g., $\mathrm{LR}=3\times 10^{-5}$), where damage accumulates across layers and results in irreversible representational collapse. Importantly, MIA AUC alone can be misleading. Models may exhibit near-complete privacy recovery while remaining functionally impaired.
Table 2: Yi-6B: MIA / F.Acc / R.Acc (%) for simple task under four unlearning settings
Phase Method LR=3$\times$$10^{-5}$, N=100 LR=5$\times$$10^{-6}$, N=100 LR=3$\times$$10^{-6}$, N=100 LR=3$\times$$10^{-5}$, N=6
MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc
Original —— 70.8 78.9 65.5 70.8 78.9 65.5 70.8 78.9 65.5 70.8 78.9 65.5
Unlearn GA 26.1 0.0 0.0 23.2 9.1 6.2 25.2 16.8 14.4 29.6 36.3 36.1
GA+GD 16.8 9.7 2.3 28.7 3.6 3.1 69.4 78.8 65.5 66.9 77.0 64.0
GA+KL 17.8 9.0 6.2 27.3 9.1 6.2 18.9 3.8 3.2 29.5 52.9 41.5
NPO 60.1 37.8 37.9 50.6 51.0 52.3 68.4 78.3 64.1 68.7 71.6 59.4
NPO+KL 59.0 64.3 55.9 65.4 77.6 64.3 66.7 78.8 65.5 67.9 67.6 56.1
RLabel 65.1 0.0 0.0 63.6 0.1 0.4 61.4 0.4 0.7 62.7 72.7 61.1
Relearn GA 74.5 2.1 1.8 68.0 80.0 65.0 68.6 80.8 65.2 68.2 70.5 58.7
GA+GD 68.1 2.2 2.6 69.8 81.2 65.1 70.0 81.8 65.5 67.0 61.6 54.4
GA+KL 70.7 1.7 1.6 68.3 81.1 64.8 70.7 81.0 63.2 65.0 66.6 56.2
NPO 70.0 57.0 45.6 68.0 82.7 65.5 69.9 81.2 65.4 68.4 71.2 59.4
NPO+KL 67.7 60.7 54.2 69.5 83.8 65.6 69.9 83.8 65.4 69.0 67.6 56.1
RLabel 69.5 4.3 2.8 70.4 80.8 65.3 70.0 80.5 65.3 65.2 72.7 61.1
Table 3: Qwen-2.5-7B: MIA / MATH / GSM8K Accuracy (%) for complex task under four settings
Phase Method LR=$3\times 10^{-5}$, $N$=6 LR=$3\times 10^{-6}$, $N$=6 LR=$5\times 10^{-6}$, $N$=6 LR=$5\times 10^{-6}$, $N$=100
MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K
Original —– 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1
Unlearn GA 5.9 0.0 0.0 0.9 0.0 0.0 3.8 0.0 0.0 5.5 0.0 0.0
NPO 95.9 0.0 0.2 97.4 21.5 74.1 67.4 24.1 71.8 94.7 0.0 0.4
RLabel 35.5 0.0 0.0 69.6 0.0 1.5 11.2 0.0 0.0 2.9 0.0 0.0
Relearn GA 97.6 0.0 1.1 99.3 5.1 83.2 99.4 9.3 77.8 99.2 0.0 0.0
NPO 95.8 0.0 0.0 99.4 4.7 82.6 99.4 16.5 75.7 99.2 0.0 0.0
RLabel 99.5 0.0 0.0 99.3 5.3 83.3 99.3 10.0 77.2 99.6 0.0 0.0
4 A Unified Representational Analysis
4.1 Representational Analysis Tools
We monitor representational drift with four layer-wise diagnostics: PCA Similarity, PCA Shift, CKA, and the diagonal Fisher Information Matrix (FIM), summarized in Figure 1(b); implementation details (including their precise definitions) are deferred to Appendix A.3.
PCA Similarity & Shift. For each layer $i$, we collect activations $\mathbf{H}_{i}^{\text{orig}}$, $\mathbf{H}_{i}^{\text{unl}}$, and $\mathbf{H}_{i}^{\text{rel}}$ on a probe set $\mathcal{X}$. Let $\mathbf{c}_{i,1}^{(*)}$ and $p_{i,1}^{(*)}$ be he first principal direction and its mean projection for state (){$(*)\!\in\!\{\text{orig},\text{unl},\text{rel}\}$. The cosine between $\mathbf{c}_{i,1}^{\text{orig}}$ and $\mathbf{c}_{i,1}^{(*)}$ yields PCA similarity; the signed difference $p_{i,1}^{(*)}-p_{i,1}^{\text{orig}}$ gives PCA shift. Small angles and shifts indicate stable features; otherwise, catastrophic forgetting [36].
Centered Kernel Alignment (CKA). With centered activation matrices $X_{i}^{\text{orig}}$ and $X_{i}^{(*)}$, we compute ${\rm CKA}(X_{i}^{\text{orig}},X_{i}^{(*)})\!\in\![0,1]$; values $\approx 1$ mean nearly identical subspaces, those $\approx 0$ are orthogonal.
Fisher information. We estimate the diagonal empirical FIM by averaging squared gradients over $\mathcal{X}$. Comparing $\mathrm{FIM}^{\text{orig}}$, $\mathrm{FIM}^{\text{unl}}$, and $\mathrm{FIM}^{\text{rel}}$ reveals how unlearning flattens the loss landscape and whether relearning restores parameter importance [14; 11].
We compute all diagnostics not only on the forget set but also on the retain set and on unrelated data, allowing us to distinguish targeted forgetting from broader representational disruption.
(a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 2: Layer-wise PCA Similarity for GA on Yi-6B (simple task). (a–c) vary LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ at $N=100$; (d–f) vary $N\in\{6,50,100\}$ at LR $=3\times 10^{-5}$. Similarity near 1 signals reversible (catastrophic) forgetting; sustained low similarity signals irreversible (catastrophic) forgetting.
(a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 3: PCA Shift for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. Shift magnitude reflects feature displacement: large, unrecovered shifts indicate severe, irreversible (catastrophic) forgetting, while small or fully recovered shifts indicate mild, reversible (catastrophic) forgetting.
4.2 Representational Results
Principal Component Analysis: Similarity and Shift. Figures 2 and 3 show that, in continual unlearning, larger learning rates or more removal requests drive PCA Similarity sharply downward and leave large, unrecovered PCA Shifts—clear evidence of irreversible catastrophic forgetting. Under gentler hyper-parameters, similarity stays high and shifts remain bounded; relearning then realigns both metrics, signalling reversible catastrophic forgetting. Remarkably, the choice of relearning data and analyzed data—forget, retain, or even unrelated—makes little difference: all three restore the feature geometry, implying the knowledge was suppressed rather than erased (Figure 10 and  14). Due to space limitations, complete figures, single-unlearning results, and additional methods appear in Appendix A.5 and Appendix A.6.
Centered Kernel Alignment Analysis. Figure 4 tracks layer-wise CKA. With gentle unlearning, CKA stays near 1 and relearning restores it completely—typical of reversible catastrophic forgetting. Stronger updates or many requests push deep-layer CKA sharply lower, and alignment cannot be fully recovered, signalling irreversible catastrophic forgetting. The choice of relearning or probe data—forget, retain, or unrelated—barely affects this outcome: once relearning begins, latent structure resurfaces regardless of input (Figure 18). Due to space limitations, complete figures and additional settings appear in Appendix A.5 and Appendix A.6.
(a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 4: CKA for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. High CKA (near 1) indicates strong subspace alignment and reversible (catastrophic) forgetting, whereas low CKA (near 0) reflects severe representational drift and irreversible (catastrophic) forgetting.
Fisher Information Analysis. Continual unlearning progressively flattens the loss landscape (Figures 5. Higher learning rates or larger request counts push the FIM spectra sharply left, and—at extreme settings—the shift persists after relearning, signalling irreversible catastrophic forgetting. Under gentler hyper-parameters the spectra recentre, indicating reversible catastrophic forgetting. Relearning with forget, retain, or unrelated data realigns the FIM almost equally well, confirming that the lost sensitivity is suppressed rather than erased (Figure 34-38). Due to space limitations, complete figures and additional settings appear in Appendix A.5 and Appendix A.6.
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) $N=6$, Layer 31
(e) $N=50$, Layer 31
(f) $N=100$, Layer 31
Figure 5: FIM (layer 31) for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. Larger leftward shifts indicate a greater flattening of the loss landscape and irreversible (catastrophic) forgetting; spectra that recenter on the original peak denote reversible (catastrophic) forgetting.
4.3 Representational Theoretical Analysis
To explain the empirical distinction between reversible and irreversible (catastrophic) forgetting, we present a perturbation model linking unlearning updates to structural collapse across layers. Consider an $L$-layer feedforward neural network $f(x)=\sigma(W_{L}\,\sigma(\cdots\sigma(W_{1}x)\cdots))$, with activations $\sigma$ and weights ${W_{i}}_{i=1}^{L}$. Unlearning is modeled as layer-wise perturbations $\widetilde{W}_{i}=W_{i}+E_{i}$ with $|E_{i}|=O(\mathrm{LR},N)$, where $\mathrm{LR}$ is the learning rate and $N$ the number of unlearning steps. A Neumann-series expansion yields $\widetilde{f}(x)-f(x)=\sum_{\emptyset\neq S\subseteq\{1,\dots,L\}}(W_{L}\circ\cdots\circ E_{i_{k}}\circ\cdots\circ W_{1})(x)$.
When small perturbations are confined to a few layers, first-order terms dominate, and the effect is reversible (catastrophic) forgetting. In contrast, comparable perturbations spread across many layers, higher-order terms accumulate, producing irreversible (catastrophic) forgetting.
PCA Similarity. Let $X_{i}$ and $Y_{i}=X_{i}+E^{\prime}_{i}$ be the centered activations at layer $i$ before and after unlearning. By Davis–Kahan theorem [5], $\cos\angle(\mathbf{c}_{i}^{\text{orig}},\mathbf{c}_{i}^{\text{upd}})\approx 1-O(\|E_{i}^{\prime}\|/(\lambda_{1,i}-\lambda_{2,i})),$ with top two eigenvalues $\lambda_{1,i},\lambda_{2,i}$. The layer-averaged PCA similarity is $\bar{S}_{\mathrm{PCA}}\approx 1-O((1/L)\sum_{i}\|E^{\prime}_{i}\|)$.
PCA Shift. Along the first principal component, the activation-centroid shift is $\Delta p_{i}=\mu_{i,1}^{\text{upd}}-\mu_{i,1}^{\text{orig}}=O(\|E_{i}^{\prime}\|)$. Large perturbations $\|E_{i}^{\prime}\|$ spanning many layers cause irreversible representational drift; otherwise, the shifts remain localized and reversible.
CKA. Let $\widetilde{K}_{Y_{i}}=\widetilde{K}_{X_{i}}+\Delta K_{i}$ be the perturbed Gram matrix. Then, $\mathrm{CKA}_{i}$ is computed as $1-O(\|\Delta K_{i}\|_{*}/\|\widetilde{K}_{X_{i}}\|_{*})$, which implies that $\bar{C}\approx 1-O((1/L)\sum_{i}\|\Delta K_{i}\|_{*})$.
Fisher Information. Given update $\delta w_{i}=O(\|E_{i}\|)$, the Fisher diagonal behaves as $F_{ii}(w+\delta w)=F_{ii}(w)+O(\|\delta w_{i}\|)$, so the average Fisher becomes $\bar{F}=(1/P)\sum_{i}F_{ii}=F_{0}-O((1/P)\sum_{i}\|E_{i}\|)$.
Token-level metrics, such as accuracy, MIA, and AUC, may report total collapse even when the model’s internal geometry is largely preserved. In reversible forgetting, a few parameter changes (e.g., in output heads or layer norms) can drastically perturb token probabilities while leaving deeper representations intact. For the soft-max output $\log p(y|x;\theta+\delta\theta)\approx\log p(y|x;\theta)+\nabla_{\theta}\log p(y|x;\theta)^{\top}\delta\theta+O(\|\delta\theta\|^{2}).$ a small $\delta\theta$ in high-sensitivity regions (near the logits) can dominate the first-order term, producing large drops in accuracy or anomalous AUC scores despite minimal representational drift.
Unlearning on the forget set applies a weight update $E_{i}=LR\times\,\nabla_{W_{i}}L(D_{f}),$ which both removes over-fitting to $\mathcal{D}_{f}$ and accentuates its principal feature subspace. After relearning, parameters may return close to their originals, yet amplified patterns persist, sometimes yielding better performance on an augmented $\mathcal{D}_{f}$ than the baseline. Therefore, unlearning can inadvertently act as a contrastive regularizer, further illustrating the mismatch between surface metrics and feature subspaces.
Summary. Reversible (catastrophic) forgetting occurs when perturbations touch only a few layers; PCA similarity/shift, CKA, and FIM remain near baseline. In contrast, irreversible (catastrophic) forgetting emerges from large, distributed updates that collapse the model’s representational structure.
Unlearning acts as a contrastive perturbation: it removes memorized content yet “reinforces” salient features of the forget set, so subsequent relearning can even outperform the original model on related inputs, revealing its “dual role” in both removal and refinement.
Token-level metrics (e.g., accuracy, MIA) are overly sensitive to small shifts in high-impact parameters and can misclassify the regime. Structural diagnostics, complemented by augmented evaluation, can provide a more reliable assessment of whether forgetting is truly irreversible.
5 Discussion and Takeaways
Beyond theoretical justifications in Section 4, we summarize main empirical and analytical insights.
(1) Single vs. continual unlearning, and the role of GA/RLabel.
Single unlearning rarely produces irreversible collapse: performance is recoverable and representational drift is slight. Continual unlearning, especially with large learning rates, often drives the model into permanent failure: $\sim 100$ sequential requests can push both forget- and retain-set accuracy to near zero. GA and RLabel already over-forget in the single scenario and magnify this damage when applied continually. Adding retain-set terms, as in GA+KL or NPO(+KL) [33; 35; 31], markedly improves stability.
(2) Collapse stems from structural drift, not true erasure.
PCA-Similarity/Shift, CKA, and the FIM consistently expose this breakdown: irreversible collapse coincides with large rotations of principal directions, centroid shifts, and vanishing Fisher mass across many layers. When perturbations remain local (small LR, few unlearning requests), these diagnostics stay near baseline—reversible forgetting. Token-level metrics alone are unreliable: small updates in high-sensitivity parameters (e.g., logits or layer norms) can tank accuracy or inflate MIA AUC while internal geometry is intact.
(3) Unlearning can act as implicit augmentation.
In several continual runs, subsequent relearning on the forget set can often yield higher accuracy than that of the original model. This surprising outcome suggests that unlearning is not merely a memory deletion mechanism but may also serve as a form of implicit contrastive regularization. As detailed in Section 4, unlearning amplifies the feature subspace associated with the forget set, and relearning on augmented inputs can reinforce semantic structure while promoting robustness. This process reorganizes internal representations to better capture generalizable patterns, acting as a form of curriculum learning.
(4) Diagnostics guide irreversible (benign) forgetting.
Tools such as PCA Similarity and Shift, CKA, and Fisher Information reveal not just the presence of structural drift, but also where and how it arises across layers. This enables targeted control: effective unlearning can be guided toward perturbing only those parameters responsible for the forget set, while preserving the representation structure on the retain set and unrelated data. This opens a path to targeted, irreversible forgetting, a permanent and isolated removal of information without collateral collapse, offering actionable insights for building safer unlearning algorithms.
6 Conclusion
We revisit machine unlearning for LLMs through a systematic study of reversibility. Token-level metrics alone can mislead: a model may appear collapsed yet remain fully recoverable.
To diagnose this gap, we introduce a feature-space toolkit—PCA Similarity, PCA Shift, CKA, and FIM—that cleanly separates reversible from irreversible catastrophic forgetting. Our empirical and theoretical results show that true forgetting arises only when many layers undergo coordinated, large-magnitude perturbations; by contrast, minor updates in high-sensitivity regions (e.g., output logits) can slash accuracy or inflate perplexity while leaving internal representations intact.
These findings call for evaluation protocols that go beyond token-level scores and for algorithms that actively control representational drift. We further observe that unlearning, followed by proper relearning, can refine representations and even boost downstream performance. Together, these insights chart a path toward safer, more interpretable unlearning in LLMs.
Appendix A Appendix
A.1 Limitations
Our experiments target two LLMs and a handful of tasks and unlearning methods; although our diagnostic framework is model-agnostic and designed to scale, empirical validation on much larger models and production-scale pipelines remains to be done. The constrained relearning protocol and selected metrics provide clear insights into representational drift but are not exhaustive and do not offer formal privacy guarantees. Extending these analyses to diverse architectures, richer unlearning strategies, and rigorous theoretical bounds is an important direction for future work.
A.2 Broader impacts
This work provides new tools and theoretical insights for evaluating and guiding machine unlearning in LLMs, with the potential to improve privacy guarantees and regulatory compliance (e.g., GDPR “right to be forgotten”). By distinguishing reversible from irreversible forgetting, our diagnostics can help developers ensure that sensitive or copyrighted material is truly expunged rather than merely hidden. At the same time, adversaries could exploit reversible forgetting pathways to mask malicious or biased content and then restore it later, highlighting the need for robust defenses. The computational overhead of continual unlearning and repeated diagnostics may increase energy consumption, underscoring the importance of efficient implementations. Finally, by exposing the fragility of token-level metrics, our work advocates for more trustworthy evaluation standards that balance privacy, utility, and environmental considerations in real-world deployments.
A.3 Detailed Analysis Tools
PCA Similarity and PCA Shift.
For each Transformer layer, we perform PCA on the hidden activations of the original and updated models. Let $\mathbf{c}_{i,1}^{\text{orig}}$ and $\mathbf{c}_{i,1}^{\text{upd}}$ denote the first principal component (PC1) directions of layer $i$. The PCA Similarity is defined as
$\operatorname{PCA\text{-}Sim}(i)=\cos\bigl(\mathbf{c}_{i,1}^{\text{orig}},\mathbf{c}_{i,1}^{\text{upd}}\bigr{missing})=\frac{(\mathbf{c}_{i,1}^{\text{orig}})^{\top}\mathbf{c}_{i,1}^{\text{upd}}}{\lVert\mathbf{c}_{i,1}^{\text{orig}}\rVert\,\lVert\mathbf{c}_{i,1}^{\text{upd}}\rVert}\in[-1,1],$
where values near $1$ indicate stable directional alignment, and values near $-1$ suggest a near-orthogonal shift in dominant directions.
To capture translational drift, we also compute the mean projection of activations along PC1 and PC2:
$\operatorname{PCA\text{-}Shift}(i)=p_{1,\text{upd}}-p_{1,\text{orig}},\qquad\operatorname{Principle}(i)=p_{2,\text{upd}},$
where $\operatorname{PCA\text{-}Shift}$ quantifies displacement along PC1 and $\operatorname{Principle}$ captures orthogonal deviation along PC2. These metrics reflect how the representation center drifts within the top subspace.
Centered Kernel Alignment (CKA).
To assess subspace alignment, we use linear Centered Kernel Alignment (CKA) [15], which compares activation matrices $X,Y\in\mathbb{R}^{N\times D}$ from before and after unlearning. First, we compute the centered Gram matrices:
$\widetilde{K}_{X}=HXX^{\top}H,\qquad\widetilde{K}_{Y}=HYY^{\top}H,\qquad H=I_{N}-\tfrac{1}{N}\mathbf{1}\mathbf{1}^{\top}.$
The CKA score is then given by:
$\operatorname{CKA}(X,Y)=\frac{\operatorname{Tr}(\widetilde{K}_{X}\widetilde{K}_{Y})}{\sqrt{\operatorname{Tr}(\widetilde{K}_{X}^{2})}\sqrt{\operatorname{Tr}(\widetilde{K}_{Y}^{2})}}\in[0,1],$
where values near $1$ indicate highly overlapping subspaces, and values near 00 signal near-orthogonality.
Fisher Information.
To measure parameter-level importance, we compute the diagonal of the empirical Fisher Information Matrix (FIM). For each parameter $w_{i}$ and input distribution $\mathcal{D}_{\text{dis}}$, the diagonal entry is approximated as:
$\operatorname{FIM}_{ii}\approx\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{\text{dis}}}\left[\left(\partial_{w_{i}}\log p(y\mid\mathbf{x};\mathbf{w})\right)^{2}\right].$
Larger values indicate that $w_{i}$ has a stronger influence on the model’s predictions. A substantial leftward shift in the Fisher spectrum after unlearning implies a flattened loss landscape and diminished parameter sensitivity.
Together, these tools form a feature-space diagnostic suite: FIM captures global sensitivity, CKA measures subspace preservation, and PCA-based metrics expose fine-grained geometric drift across layers—enabling a robust assessment of representational degradation during unlearning.
A.4 Complete Experiment
Table 4: Yi-6B simple-task metrics under four $(\mathrm{LR},N)$ settings. For each block: forget/retain perplexity (F.Ppl / R.Ppl), forget/retain accuracy (F.Acc / R.Acc), CommonsenseQA (CSQA), GSM8K, and membership-inference AUC (MIA).
Phase Method F.Ppl R.Ppl F.Acc R.Acc CSQA GSM8K MIA
LR=3×$1003\times 10^{-5},\,N=$ 100
Original 3.8 7.8 78.9 65.5 73.1 39.6 70.9
Unlearn GA $\infty$ $\infty$ 0.0 0.0 19.3 0.0 26.1
GA+GD $\infty$ $\infty$ 9.7 2.3 19.7 0.0 16.8
GA+KL $\infty$ $\infty$ 9.0 6.2 19.6 0.0 17.8
NPO 31296.5 597.9 37.8 37.9 62.2 1.0 60.1
NPO+KL 348080.2 4482.0 64.3 55.9 64.9 1.4 59.0
Rlable 63791.7 65903.4 0.0 0.0 20.9 0.0 65.1
Relearn GA 137094.5 758443.5 2.1 1.8 19.7 0.0 74.5
GA+GD 5274.5 9568.6 2.2 2.6 19.6 0.0 68.1
GA+KL 5037.1 15019.9 1.7 1.6 20.6 0.0 70.7
NPO 16.6 41.7 57.0 45.6 51.8 0.6 70.0
NPO+KL 21.8 16.2 60.7 54.3 48.0 0.9 67.7
Rlable 4056.1 15048.6 4.3 2.8 19.7 0.0 69.5
LR=5×$1005\times 10^{-6},\,N=$ 100
Unlearn GA $\infty$ $\infty$ 9.1 6.2 19.6 0.0 23.2
GA+GD $\infty$ $\infty$ 3.6 3.1 24.5 0.0 28.7
GA+KL $\infty$ $\infty$ 9.1 6.2 19.6 0.0 27.3
NPO 3017.7 1110.6 50.1 52.3 72.9 37.5 50.6
NPO+KL 38.5 232.4 77.6 64.3 73.1 37.6 65.4
Rlable 57035.4 53377.1 0.1 0.4 19.1 0.0 63.6
Relearn GA 3.7 7.8 80.0 64.9 70.2 39.9 68.0
GA+GD 3.6 7.6 81.2 65.1 72.1 39.0 69.8
GA+KL 3.6 8.4 81.1 64.8 71.6 40.7 68.3
NPO 3.5 7.6 82.7 65.5 74.0 39.7 68.0
NPO+KL 3.5 7.8 83.8 65.6 74.1 39.7 69.5
Rlable 3.6 7.7 80.8 65.3 71.8 39.2 70.3
LR=3×$1003\times 10^{-6},\,N=$ 100
Unlearn GA $\infty$ $\infty$ 16.8 14.4 69.5 12.3 25.2
GA+GD 3.3 7.6 78.8 65.5 77.0 37.5 69.4
GA+KL $\infty$ $\infty$ 35.4 40.6 63.2 18.3 18.9
NPO 3.7 7.9 78.3 65.0 73.3 38.7 68.4
NPO+KL 3.8 8.1 78.4 65.1 73.6 38.6 66.7
Rlable 36794.7 32562.0 3.8 3.2 19.3 2.2 61.4
Relearn GA 3.7 7.6 80.8 65.2 73.4 39.9 68.6
GA+GD 3.6 7.4 81.8 65.5 72.1 39.0 70.0
GA+KL 3.6 10.3 81.0 63.3 67.2 40.7 70.7
NPO 3.5 7.5 81.2 65.4 72.9 39.7 69.9
NPO+KL 3.5 7.5 83.8 65.5 73.0 39.7 69.9
Rlable 3.6 7.6 80.5 65.3 72.2 39.2 70.0
LR=$3\times 10^{-5},\,N=6$
Unlearn GA inf inf 36.3 36.1 69.1 5.8 29.6
GA+GD 209.3 20.6 77.0 64.0 70.0 37.8 66.9
GA+KL inf inf 53.0 41.5 68.3 2.0 29.5
NPO 12.3 10.7 71.6 59.4 71.7 24.7 68.7
NPO+KL 8.9 10.7 74.7 62.1 72.8 32.2 67.9
Rlable 51589.2 40622.9 0.4 0.7 19.8 0.0 62.6
Relearn GA 6.8 11.4 70.5 58.7 64.5 18.4 68.2
GA+GD 12.3 11.5 61.6 54.4 61.3 7.3 67.1
GA+KL 17.1 11.6 66.6 56.2 60.6 3.0 65.0
NPO 6.0 11.6 71.2 59.4 59.4 2.0 68.4
NPO+KL 7.3 11.6 67.6 56.1 42.9 1.6 69.0
Rlable 6.4 11.4 72.7 61.1 67.5 28.9 65.2
Table 4 summarizes token-level performance under four unlearning configurations on Yi-6B, varying both learning rates and request counts. Most metrics—including forget/retain accuracy, perplexity, CSQA, and GSM8K—follow consistent trends: milder methods (e.g., NPO, NPO+KL) yield moderate degradation and support effective recovery via relearning, while aggressive methods (e.g., GA, RL) under high learning rates lead to severe, often irrecoverable performance drops. These results align closely with our theoretical distinction: reversible (catastrophic) forgetting corresponds to localized perturbations and structurally recoverable models, while irreversible (catastrophic) forgetting reflects widespread, irreversible damage.
MIA, however, behaves differently. Even when downstream performance collapses, MIA scores often remain high and recover rapidly. This suggests MIA is especially sensitive to shallow changes (e.g., output logits or normalization), rather than deeper representational shifts. As a result, MIA often correlates with reversible forgetting—capturing surface-level instability without revealing structural collapse. These observations reinforce that while token-level metrics offer partial signals, only internal diagnostics can reliably differentiate reversible from irreversible forgetting.
A.5 Single Unlearning
Figure 6 illustrates feature-level changes under single unlearning using PCA Similarity, PCA Shift, CKA, and Fisher Information. In (a), PCA Similarity remains consistently high across layers, with cosine scores near 1, indicating that dominant activation directions are well preserved. Slight dips in shallow and final layers are quickly restored after relearning, suggesting minimal and reversible geometric drift. Subfigure (b) confirms that PC1 shifts and orthogonal deviations are small, with relearned centers closely matching the original. In (c), CKA shows near-perfect alignment between the original, unlearned, and relearned representations, reinforcing the conclusion that subspace structure remains intact. Fisher Information spectra in (d–f) reveal only mild leftward shifts, indicating slight loss flattening and reduced parameter sensitivity, which are fully recovered after relearning. Overall, these results confirm that single unlearning causes only minor, reversible structural perturbations—highlighting the fragility of token-level evaluations in capturing irreversible forgetting.
(a) Simple (Single Unlearning)
(b) Simple (Single Unlearning)
(c) Simple (Single Unlearning)
(d) Simple (Single Unlearning)
(e) Simple (Single Unlearning)
(f) Simple (Single Unlearning)
Figure 6: Single unlearning analysis on Yi-6B for GA under a simple task. PCA Similarity, PCA Shift, CKA, and Fisher information (FIM) distributions are reported across key layers to assess representation and parameter changes before and after unlearning and relearning.
A.6 Detailed Analysis Results
(a) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(k) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(n) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(h) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(k) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(n) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 8: PCA Similarity Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
[…]
(a) Simple (Relearned by forget set)
(b) Simple (Relearned by Retain set)
(c) Simple (Relearned by unrelated data)
(d) Simple (input data = forget set)
(e) Simple (input data = retain set)
(f) Simple (input data = unrelated data)
Figure 10: PCA Similarity Analysis for GA under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): PCA similarity is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(k) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(n) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(h) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(k) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(n) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 12: PCA Shift Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple (Relearned by forget set)
(b) Simple (Relearned by retain set)
(c) Simple (Relearned by unrelated data)
(d) Simple (input data = forget set)
(e) Simple (input data = retain set)
(f) Simple (input data = unrelated data)
Figure 14: PCA Shift Analysis under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): PCA shift is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(k) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(l) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(n) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(o) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(h) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(i) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(k) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(l) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(n) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(o) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 16: CKA Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(b) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(c) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(d) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(e) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(f) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(g) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(h) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(i) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
[…]
(a) Simple (Relearned by forget set)
(b) Simple (Relearned by retain set)
(c) Simple (Relearned by unrelated data)
(d) Simple (input data = forget set)
(e) Simple (input data = retain set)
(f) Simple (input data = unrelated data)
Figure 18: CKA Analysis under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): CKA is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 22
(e) LR=$5\times 10^{-6}$, Layer 22
(f) LR=$3\times 10^{-5}$, Layer 22
(g) LR=$3\times 10^{-6}$, Layer 13
(h) LR=$5\times 10^{-6}$, Layer 13
(i) LR=$3\times 10^{-5}$, Layer 13
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) $N=6$, Layer 28
(b) $N=50$, Layer 28
(c) $N=100$, Layer 28
(d) $N=6$, Layer 22
(e) $N=50$, Layer 22
(f) $N=100$, Layer 22
(g) $N=6$, Layer 13
(h) $N=50$, Layer 13
(i) $N=100$, Layer 13
(j) $N=6$, Layer 4
(k) $N=50$, Layer 4
(l) $N=100$, Layer 4
(m) $N=6$, Layer 1
(n) $N=50$, Layer 1
(o) $N=100$, Layer 1
Figure 25: FIM for GA Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 26: FIM for GA+GD Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 27: FIM for GA+KL Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 28: FIM for NPO Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 29: FIM for NPO+KL Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 30: FIM for Rlable Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) Simple (Relearned by forget set), Layer 31
(b) Simple (Relearned by retain set), Layer 31
(c) Simple (Relearned by unrelated data), Layer 31
(d) Simple (input data = forget set), Layer 31
(e) Simple (input data = retain set), Layer 31
(f) Simple (input data = unrelated data), Layer 31
Figure 34: FIM in layer 31 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple (Relearned by forget set), Layer 25
(b) Simple (Relearned by retain set), Layer 25
(c) Simple (Relearned by unrelated data), Layer 25
(d) Simple (input data = forget set), Layer 25
(e) Simple (input data = retain set), Layer 25
(f) Simple (input data = unrelated data), Layer 25
Figure 35: FIM in layer 25 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple (Relearned by forget set), Layer 16
(b) Simple (Relearned by retain set), Layer 16
(c) Simple (Relearned by unrelated data), Layer 16
(d) Simple (input data = forget set), Layer 16
(e) Simple (input data = retain set), Layer 16
(f) Simple (input data = unrelated data), Layer 16
Figure 36: FIM in layer 16 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple (Relearned by forget set), Layer 4
(b) Simple (Relearned by retain set), Layer 4
(c) Simple (Relearned by unrelated data), Layer 4
(d) Simple (input data = forget set Layer), 4
(e) Simple (input data = retain set Layer), 4
(f) Simple (input data = unrelated data), Layer 4
Figure 37: FIM in layer 4 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple (Relearned by forget set), Layer 1
(b) Simple (Relearned by retain set), Layer 1
(c) Simple (Relearned by unrelated data), Layer 1
(d) Simple (input data = forget set), Layer 1
(e) Simple (input data = retain set), Layer 1
(f) Simple (input data = unrelated data), Layer 1
Figure 38: FIM in layer 1 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
A.6.1 Principal Component Analysis: Similarity and Shift
Across the same hyper-parameter grid, Figures 7– 13 plot the PCA–Shift trajectories—layer-wise displacements of activation centroids along the first two principal directions. For GA-based objectives the pattern mirrors their Similarity curves: as LR rises the orange triangles (unlearned) shoot far from the blue circles (original), especially in deeper layers, and the green squares (relearned) fail to return, producing long grey rays that diagnose irreversible drift. With GA+GD the spread is smaller but still widens sharply at $3{\times}10^{-5}$, confirming that doubling the loss term does not prevent global collapse.
NPO and NPO+KL behave differently. Even at aggressive LR the shifts remain tightly clustered—most layers move $<\!100$ units on pc1 and almost none on pc2—and green squares consistently fall back a little onto the original line segment. RLabel shows an intermediate picture: early layers barely move, while late layers fan out as LR or $N$ grow; nonetheless the rays shorten markedly after relearning, indicating that most distortion is still recoverable.
Task complexity amplifies divergence. On Qwen-2.5-7B the GA rays explode more quickly (Figure 13c,f,i), spanning thousands of pc1 units and driving pc2 to extreme negative values—the hallmark of a high-order, multi-layer perturbation predicted by our theory in Section 4. Conversely, NPO’s clusters expand only modestly, even under the same LR, and contract once relearning is applied.
Taken together, PCA-Shift complements PCA-Similarity: Similarity captures angular mis-alignment, while Shift quantifies translational drift. Their joint reading confirms that GA (with or without GD/KL) is prone to large, irreversible representational displacements, whereas NPO variants and, to a lesser extent, RLabel confine shifts to a regime that remains correctable—consistent with the reversible versus irreversible forgetting boundary observed in our utility experiments.
A.6.2 Centered Kernel Alignment Analysis
Figures 15– 17 report layer-wise linear CKA between the original model and its unlearned / relearned counterparts. Across both Yi-6B and Qwen-2.5-7B, GA again stands out: as LR or $N$ grows its CKA curve close to zero in the final third of the network and never returns, revealing a deep sub-space fracture that matches the irreversible PCA trends. GA+GD and GA+KL modestly attenuate this dip, but still fail to restore full alignment after relearning.
By contrast, NPO and NPO+KL keep CKA higher GA through almost all layers—even under $3{\times}10^{-5}$ or $N{=}100$—and relearning lifts the few layers back to baseline, confirming their perturbations are lower than GA series. RLabel occupies an intermediate position: as LR or N increases, its CKA curve drops rapidly—mirroring GA’s behavior—and ultimately exhibits irreversible forgetting.
Task complexity does not change the ordering but amplifies the gaps: on the math-heavy Qwen benchmark GA’s tail layers fall to almost zero at high LR, whereas NPO keeps higher than GA. Taken together with the PCA-Shift results, CKA shows that only GA-style objectives consistently destroy the encoder–decoder sub-space, while NPO families maintain higher stability than GA series and RLabel induces a moderate, recoverable tilt.
A.6.3 Fisher Information Analysis
[…]

### v2

@@
 Abstract
-Unlearning in large language models (LLMs) is intended to remove the influence of specific data, yet current evaluations rely heavily on token-level metrics such as accuracy and perplexity.
-We show that these metrics can be misleading: models often appear to forget, but their original behavior can be rapidly restored with minimal fine-tuning, revealing that unlearning may obscure information rather than erase it.
-To diagnose this phenomenon, we introduce a representation-level evaluation framework using PCA-based similarity and shift, centered kernel alignment, and Fisher information.
-Applying this toolkit across six unlearning methods, three domains (text, code, math), and two open-source LLMs, we uncover a critical distinction between reversible and irreversible forgetting.
-In reversible cases, models suffer token-level collapse yet retain latent features; in irreversible cases, deeper representational damage occurs.
-We further provide a theoretical account linking shallow weight perturbations near output layers to misleading unlearning signals, and show that reversibility is modulated by task type and hyperparameters.
-Our findings reveal a fundamental gap in current evaluation practices and establish a new diagnostic foundation for trustworthy unlearning in LLMs.
-We provide a unified toolkit for analyzing LLM representation changes under unlearning and relearning: https://github.com/XiaoyuXU1/Representational_Analysis_Tools.git.
+Unlearning in large language models (LLMs) aims to remove specified data, but its efficacy is typically assessed with task-level metrics like accuracy and perplexity.
+We demonstrate that these metrics are often misleading, as models can appear to forget while their original behavior is easily restored through minimal fine-tuning.
+This phenomenon of reversibility suggests that information is merely suppressed, not genuinely erased.
+To address this critical evaluation gap, we introduce a representation-level analysis framework.
+Our toolkit comprises PCA-based similarity and shift, centered kernel alignment (CKA), and Fisher information, complemented by a summary metric, the mean PCA distance, to measure representational drift.
+Applying this framework across six unlearning methods, three data domains, and two LLMs, we identify four distinct forgetting regimes based on their reversibility and catastrophicity.
+Our analysis reveals that achieving the ideal state–irreversible, non-catastrophic forgetting–is exceptionally challenging.
+By probing the limits of unlearning, we identify a case of seemingly irreversible, targeted forgetting, offering new insights for designing more robust erasure algorithms.
+Our findings expose a fundamental gap in current evaluation practices and establish a representation-level foundation for trustworthy unlearning.
+[^1] https://github.com/XiaoyuXU1/Representational_Analysis_Tools
 1 Introduction
 Large language models (LLMs), trained on massive corpora, have achieved remarkable success across diverse tasks, yet their capacity to memorize training snippets poses acute ethical, legal, and security risks.
-Memorization can unintentionally disclose sensitive, harmful, or copyrighted text [25; 13; 30], conflicting with emerging regulations, such as the EU’s Right to be Forgotten [9].
-Machine unlearning seeks to mitigate this threat by making a model act as though specified data were never seen [2].
-Numerous methods have been proposed for LLMs [33; 12; 6; 26; 18; 17; 19; 20; 31], their success usually judged by token-level metrics, such as accuracy or perplexity.
-However, a pivotal question remains largely unexplored: Does LLM unlearning truly erase information, or merely suppress it, poised to “resurface” at the slightest nudge?
-Empirically, many unlearning methods may just appear effective: after unlearning, a model shows near-zero accuracy or high perplexity on the forget set, but a brief fine-tuning step (even on unrelated data) can quickly restore its original behaviour (see Figure 1).
-This exposes a huge gap between surface-level metrics and the model’s internal state [21; 23], casting doubt on compliance and safety claims.
-If information is recoverable via simple fine-tuning (or relearning), can we truly claim that it has been “forgotten?” What looks like memory loss may, in fact, be a shallow perturbation.
-In this work, we perform the first systematic analysis of reversibility of LLM unlearning, covering both single-shot and continual settings.
-The continual setting allows multiple unlearning requests over time, which we believe is a more prominent and realistic setting where the deployment of such unlearning strategies are often facing a dynamic environment.
-We show that standard token-level metrics prove insufficient—they can collapse even when the underlying representations remain intact.
-To probe deeper, we introduce a diagnostic toolkit for representational analysis, featuring PCA subspace similarity and shift [36], centered kernel alignment (CKA) [15], and Fisher information [3].
-Our toolkit uncovers two distinct regimes of unlearning:
-1) Reversible (catastrophic) forgetting: performance collapses, but feature subspaces are largely preserved, enabling rapid recovery, and 2) Irreversible (catastrophic) forgetting: collapse coincides with substantial representational drift, making recovery difficult/impossible.
-Surprisingly, both yield similar results under token-level metrics, underscoring the need for deeper representational analysis.
-We further show that the transition between reversible and irreversible forgetting depends not only on the volume of unlearning requests but also on hyperparameters such as learning rate.
-Modest weight perturbations—especially near the output layer—can lead to token-level distortions without altering feature geometry, making “forgotten” knowledge easily recoverable.
-Hence, evaluating the effectiveness of LLM unlearning must go beyond superficial token-level metrics (e.g., forget accuracy declines).
-In safety- and privacy-critical settings, unlearning should be judged by its ability to achieve genuine erasure rather than simply representational collapse.
-Figure 1: (a) Token-level accuracy and CKA subspaces of reversible (top) vs. irreversible (bottom) catastrophic forgetting due to continual unlearning then relearning, (b) Our four diagnostic tools
+Memorization can unintentionally disclose sensitive, harmful, or copyrighted text [28; 14; 35], conflicting with emerging regulations, such as the EU’s Right to be Forgotten [8].
+Machine unlearning seeks to address this challenge by algorithmically erasing the influence of specified data, making a model behave as if it had never been trained on that data [2].
+While numerous unlearning methods have been developed for LLMs [40; 12; 29; 20; 19; 22; 37], their efficacy is typically assessed using task-level metrics, such as accuracy on a held-out “forget set.”
+However, these evaluations overlook a pivotal question: Does LLM unlearning achieve genuine erasure, or does it merely suppress information, leaving it poised to resurface?
+Emerging evidence suggests that many unlearning methods may only be superficially effective.
+For instance, after unlearning, a model may exhibit poor performance on the forget set, yet this “forgotten” knowledge can often be rapidly restored through a brief fine-tuning step, even on unrelated data [24; 26] (see Figure 1).
+Similarly, low-bit quantization of model weights has been shown to recover forget-set performance without any retraining, offering a simple probe for this reversibility [44].
+These findings reveal a critical gap between task-level metrics and the model’s internal state.
+If purportedly erased information can be easily recovered, the unlearning process may amount to little more than a shallow perturbation, failing to provide meaningful safety or compliance guarantees.
+While prior work has noted the phenomenon of reversibility [44], it has not been systematically investigated, particularly in the context of continual unlearning, a more realistic scenario where deletion requests arrive sequentially over a model’s lifecycle.
+This paper presents the first systematic analysis of the reversibility of LLM unlearning.
+We argue that standard task-level metrics are insufficient, as performance can collapse (a.k.a.
+catastrophic forgetting) even when the underlying feature representations remain largely intact.
+To move beyond superficial evaluations, we introduce a diagnostic toolkit for representational analysis, which includes PCA subspace similarity and shift [45], centered kernel alignment (CKA) [16], and Fisher information (FIM) [3].
+Building on this, we propose the mean PCA distance as a quantitative measure to summarize the degree of representational drift.
+With our toolkit, we build a taxonomy characterizing unlearning along two axes: reversibility and catastrophicity (collateral damage to retained knowledge).
+This allows us to distinguish four regimes: 1) Reversible, Catastrophic: Global performance collapse that is fully recoverable via relearning.
+2) Reversible, Non-Catastrophic: Targeted performance modest degradation that is easily restored.
+3) Irreversible, Catastrophic: Permanent and unrecoverable global performance collapse.
+4) Irreversible, Non-Catastrophic: Ideally, permanent erasure of target data without collateral damage.
+Crucially, we find that task-level metrics (e.g., forget accuracy) alone cannot distinguish between reversible and irreversible (catastrophic) forgetting, underscoring the necessity of our representation-level analysis; the evaluation must transcend task-level performance.
+We further show that transitions between these regimes depend on the unlearning method, the volume of unlearning requests, and key hyperparameters.
+In safety-critical applications, the central criterion for success should be the demonstrable achievement of irreversible erasure, not merely superficial performance degradation.
+Figure 1: (a) task-level accuracy and CKA subspaces of reversible (top) vs. irreversible (bottom) catastrophic forgetting due to continual unlearning then relearning, (b) Our four diagnostic tools
 Contributions. We summarize our main contributions as follows:
-  • We present the first systematic study of reversibility in both single and continual LLM unlearning, using a feature-space toolkit, including PCA similarity, PCA shift, CKA, and Fisher information.
-Our analysis distinguishes reversible from irreversible (catastrophic) forgetting.
-  • We conduct extensive experiments with six unlearning methods (GA [33], NPO [35], and RLabel, with their variants) across three datasets (arXiv papers, GitHub code [33], and NuminaMath-1.5 [16]) on Yi-6B [34] and Qwen-2.5-7B [32].
-Our results show that standard token-level metrics (e.g., accuracy, perplexity, MIA susceptibility [27]) fail to capture true forgetting behavior.
+  • We present the first systematic study of reversibility in both single and continual LLM unlearning.
+We introduce a representation-level diagnostic toolkit and a quantitative metric, the mean PCA distance, to analyze representational drift and distinguish four regimes of forgetting.
+  • We conduct extensive experiments with six unlearning methods on two LLMs (Yi-6B and Qwen-2.5-7B) across three distinct domains.
+Our results demonstrate that standard task-level metrics (e.g., accuracy, perplexity, MIA susceptibility) are insufficient for assessing the true extent of unlearning.
   • We theoretically analyze weight perturbations to explain how widespread vs.
 localized parameter changes relate to (ir)reversible forgetting.
-Small perturbations near the logits can distort token-level metrics despite intact features, hence leading to misleading assessments.
-  • Based on our findings and extra preliminary results, we propose several future directions, including using unlearning as a complementary form of data augmentation, and designing more robust unlearning algorithms that achieve genuine forgetting while avoiding representational collapse.
-2 Preliminaries
-LLM unlearning seeks to enhance privacy, improve safety, and reduce bias [33; 12; 26; 18; 17; 19].
-Most work adopts the single-unlearning paradigm: given a training corpus $\mathcal{D}$ and a designated forget set $\mathcal{D}_{f}\subseteq\mathcal{D}$, a model $\mathcal{M}$ is first trained on $\mathcal{D}$ with algorithm $\mathcal{A}$.
-An unlearning procedure $\mathcal{U}$ then transforms $\mathcal{M}$ into an unlearned model $\mathcal{M}_{f}$ that should behave as if it had never encountered $\mathcal{D}_{f}$.
-Ideally, $\mathcal{U}$ should produce a model statistically indistinguishable from one retrained on the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$: $\mathcal{M}_{f}=\mathcal{U}\!\bigl{(}\mathcal{M},\mathcal{D}_{f}\bigr{)}\approx\mathcal{M}_{r}=\mathcal{A}\!\bigl{(}\mathcal{M},\mathcal{D}_{r}\bigr{)}$.
-While current methods strike a good forget-utility balance in controlled settings [1; 8] (e.g., a fixed $\mathcal{D}_{f}$ or just a single removal request), they seldom address the practical need for continual unlearning, i.e., data owners can submit removal requests sequentially over time [1].
-Let the successive forget sets be $\mathcal{D}_{f}^{(1)},\mathcal{D}_{f}^{(2)},\ldots,\mathcal{D}_{f}^{(t)}$ (whose union is $\mathcal{D}_{f}$); the retain set after $t$ rounds is $\mathcal{D}_{r}^{(t)}$.
-The model is then updated recursively: $\mathcal{M}_{f}^{(t+1)}=\mathcal{U}(\mathcal{M}_{f}^{(t)},\mathcal{D}_{f}^{(t+1)})$, which should be similar to $\mathcal{M}_{r}=\mathcal{A}\!\bigl{(}\mathcal{M},\mathcal{D}_{r}^{(t+1)}\bigr{)}$ at any time $t$.
-Retraining LLMs is prohibitively costly, so most studies rely on empirical proxies rather than formal statistically-indistinguishable guarantees [6; 24; 18; 7].
-Evaluations track forget quality on the forget set and utility on the retain set, aiming to preserve both [24].
-While single unlearning often yields modest declines, it is fragile: brief fine-tuning—even on benign, unrelated data—swiftly revives the “forgotten” knowledge [1; 23; 21].
-The issue worsens under continual unlearning, where each round begins from an already degraded model, ultimately triggering catastrophic forgetting—a wholesale performance collapse [1; 28].
-Prior work notes this risk but does not examine its root causes.
-We hypothesize that collapse does not imply true erasure; the knowledge may remain latent in the feature space. This insight leads us to distinguish two regimes of (catastrophic) forgetting:
-Definition 1 (Reversible (Catastrophic) Forgetting).
-Let $\theta_{0}$ denote the initial model parameters, $\mathcal{D}_{f}$ the forget set, and $\mathcal{T}$ an evaluation task with metric $E(\cdot,\mathcal{T})$.
-Unlearning $\mathcal{D}_{f}$ transforms the model to $\theta_{u}$.
-If subsequent relearning on $\mathcal{D}_{f}$ (or an equivalent reconstruction set) produces parameters $\theta_{r}$ s.t.
-$E(\theta_{u},\mathcal{T})\;\ll\;E(\theta_{r},\mathcal{T})\;\approx\;E(\theta_{0},\mathcal{T}),$
-the temporary performance collapse is fully reversible; we call it reversible catastrophic forgetting.
-When the initial degradation is modest (e.g., single unlearning), we simply call it reversible forgetting.
-Definition 2 (Irreversible (Catastrophic) Forgetting).
-Using the notations of Definition 1, if
-$E(\theta_{u},\mathcal{T})\;\approx\;E(\theta_{r},\mathcal{T})\;\ll\;E(\theta_{0},\mathcal{T}),$
-we observe irreversible catastrophic forgetting: the collapse (i.e., weight perturbation) is irreversible.
-We refine this to irreversible forgetting, which further requires that the irreversible degradation be restricted to the forget set; performance on the retain set and on unrelated data must remain near their original levels.
-This condition distinguishes targeted erasure from global model failure or collapse.
-To distinguish our setting from a full retrain model, where reversibility is trivially achievable, we introduce the following restriction on the relearning phase.
+Small perturbations near the logits can distort task-level metrics despite intact features, hence leading to misleading assessments.
+  • We identify a case of seemingly irreversible, non-catastrophic forgetting, offering insights for designing more robust unlearning algorithms.
+We also highlight the potential for unlearning to serve as a form of data augmentation, improving model representations upon relearning.
+2 Preliminaries and Our Formulation
+LLM unlearning aims to remove the influence of specific data from a trained model to enhance privacy, improve safety, or mitigate bias [40; 12; 29; 20; 19; 22].
+The standard paradigm involves a training corpus $\mathcal{D}$, from which $\mathcal{D}_{f}\subseteq\mathcal{D}$ is designated as the forget set.
+A model $\mathcal{M}$ is first trained on $\mathcal{D}$ via an algorithm $\mathcal{A}$.
+An unlearning procedure $\mathcal{U}$ then transforms $\mathcal{M}$ into unlearned $\mathcal{M}_{f}$, which should ideally behave as if it were trained only on the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$.
+Formally, the goal is to statistically approximate the retrained model $\mathcal{M}_{r}$: $\mathcal{M}_{f}=\mathcal{U}(\mathcal{M},\mathcal{D}_{f})\approx\mathcal{M}_{r}=\mathcal{A}(\mathcal{D}_{r}).$
+Retraining LLMs is prohibitively costly, so most studies rely on empirical proxies rather than formal statistically-indistinguishable guarantees [27; 20; 6].
+Evaluations track forget quality on the forget set, utility, and downstream task accuracy on the retain set, aiming to preserve both dimensions.
+While current methods can achieve reasonable balances between forgetting and utility in single-shot scenarios [1; 7], they often falter in the practical continual setting, where removal requests arrive sequentially over time [1].
+For a sequence of forget sets $\mathcal{D}_{f}^{(1)},\mathcal{D}_{f}^{(2)},\ldots,\mathcal{D}_{f}^{(t)}$ (the union is $\mathcal{D}_{f}$), the retain set is $\mathcal{D}_{r}^{(t)}$ after $t$ rounds.
+The model is then updated recursively: $\mathcal{M}_{f}^{(t)}=\mathcal{U}(\mathcal{M}_{f}^{(t-1)},\mathcal{D}_{f}^{(t)})$, which should be similar to $\mathcal{M}_{r}=\mathcal{A}\!\bigl(\mathcal{M},\mathcal{D}_{r}^{(t)}\bigr),\forall t$.
+However, empirically, it often leads to catastrophic forgetting–a severe decline in performance on both forgotten and retained knowledge [1; 32].
+Single-shot unlearning is “fragile:” fine-tuning, even on benign, unrelated data, can rapidly restore the supposedly “forgotten” knowledge [1; 26; 24].
+Such fragility persists in continual unlearning as well.
+Prior work has noted this phenomenon but has not deeply investigated its underlying mechanics.
+2.1 A Taxonomy of Forgetting Regimes
+We hypothesize that this performance collapse does not necessarily equate to true information erasure; the knowledge might merely become latent or suppressed.
+To formalize this hypothesis, we introduce a taxonomy of forgetting based on two axes: catastrophicity (the extent of collateral damage to retained knowledge) and reversibility (whether forgotten knowledge can be recovered).
+Let $\theta_{0}$ be the initial model parameters, $\theta_{u}$ be the parameters after unlearning, and $\theta_{r}$ be the parameters after a subsequent relearning phase (defined below).
+We use $E(\theta,\mathcal{T})$ to denote a performance metric (e.g., accuracy) evaluated on a task set $\mathcal{T}$, which can be partitioned into a forget-related task $\mathcal{T}_{f}$ and a retain-related task $\mathcal{T}_{r}$.
+We define four distinct regimes of forgetting, summarized in Table 1.
+Definition 1 (Four Regimes of Forgetting).
+Let $\Delta_{u}(\mathcal{T})=E(\theta_{0},\mathcal{T})-E(\theta_{u},\mathcal{T})$ be the performance drop after unlearning, and $\Delta_{r}(\mathcal{T})=E(\theta_{0},\mathcal{T})-E(\theta_{r},\mathcal{T})$ be the residual drop after relearning.
+The nature of forgetting is determined by these drops on the forget set ($\mathcal{T}_{f}$) and retain set ($\mathcal{T}_{r}$).
+Catastrophic vs.
+Non-Catastrophic: Forgetting is catastrophic if utility on the retain set degrades significantly (both $\Delta_{u}(\mathcal{T}_{r})\ and\ \Delta_{u}(\mathcal{T}_{u})\gg 0$) and non-catastrophic otherwise ($\Delta_{u}(\mathcal{T}_{r})\approx 0$).
+Reversible vs.
+Irreversible: Forgetting is reversible if relearning almost recovers initial performance on forget set ($\Delta_{r}(\mathcal{T}_{f})\approx 0$) and irreversible if a significant performance drop persists ($\Delta_{r}(\mathcal{T}_{f})\gg 0$).
+The combination of these two properties yields four regimes, among which the irreversible, non-catastrophic forgetting is deemed ideal, but remains challenging to achieve in practice.
 Relearning Restriction.
-After unlearning, we briefly fine-tune $\theta_{u}$ on a small relearning set—either the cumulative forget set $\mathcal{D}_{f}=\bigcup_{t}\mathcal{D}_{f}^{(t)}$, its a distribution-similar retain set $\mathcal{D}_{r}^{(t)}$, or an unrelated out-of-distribution corpus—to obtain $\theta_{r}$ without ever revisiting the full pre-training data.
-3 Token-Level Evaluation
+To test reversibility, we must define a constrained relearning process that is distinct from a full retrain.
+After unlearning yields $\theta_{u}$, we obtain $\theta_{r}$ by briefly fine-tuning $\theta_{u}$ on a small relearning set, without access to raw pre-training data.
+This set is size-matched to $\mathcal{D}_{f}$ (i.e., total budget is $|\mathcal{D}_{f}|$) and drawn from one of three sources: (i) the forget set $\mathcal{D}_{f}$ itself, representing a worst-case recovery scenario, (ii) a retain subset $\mathcal{D}_{r}^{(t)}$ from domains similar to $\mathcal{D}_{f}$, or (iii) general-domain, out-of-distribution data.
+This ensures a fair and direct assessment of knowledge recoverability.
+[…]
+Regime Observed Description
+[…]
+[…]
+[…]
+[…]
+3 Classic (Task-Level) Evaluation Can Be Deceptive
 3.1 Experiment setup
 Models and Datasets.
-We conduct experiments on two open-source models, Yi-6B [34] and Qwen-2.5-7B [32].
-To ensure the generality of our findings, we use two dataset types: i) simple tasks—arXiv abstracts and GitHub code from [33], and ii) a complex task—NuminaMath-1.5, a recent benchmark for mathematical reasoning [16].
-All experiments were run on NVIDIA H100 GPUs.
-Unlearning algorithms We compare six canonical methods grouped into three families.
-1) Gradient-Ascent (GA) family.
-The unified goal is $\mathcal{L}\;=\;\mathcal{L}_{\text{forget}}\!\bigl{(}\mathcal{D}_{f}\bigr{)}+\;\lambda\,\mathcal{L}_{\text{retain}}\!\bigl{(}\mathcal{D}_{r}\bigr{)},$ where $\mathcal{L}_{\mathrm{forget}}$ maximizes the loss on the forget set via GA, $\mathcal{L}_{\mathrm{retain}}$ (optional) preserves utility on the retain set, and $\lambda>0$ balances the two.
-Choices for $\mathcal{L}_{\mathrm{retain}}$ give three variants: i) GA ($\mathcal{L}_{\mathrm{retain}}=0$), ii) GA+GD (standard cross-entropy on $\mathcal{D}{r}$), and iii) GA+KL (KL divergence to the reference model on $\mathcal{D}{r}$[33].
-2) Negative Preference Optimization (NPO) family.
-GA is replaced by an NPO loss that penalizes agreement with the forget set [35]: $\mathcal{L}\;=\;\mathcal{L}_{\text{NPO}}\!\bigl{(}\mathcal{D}_{f}\bigr{)}+\;\lambda\,\mathcal{L}_{\text{retain}}\!\bigl{(}\mathcal{D}_{r}\bigr{)},$ Variants mirror those above: NPO ($\mathcal{L}_{\mathrm{retain}}=0$) and NPO+KL (retain-set KL regularization).
-3) Random Label (RLabel).
-To mimic a model that never saw $\mathcal{D}_{f}$, true labels are replaced with random ones: $\mathcal{L}\;=\;\mathcal{L}_{\text{RLabel}}\!\bigl{(}\mathcal{D}_{f}\bigr{)},$ inducing near-uniform predictions without GA/negative rewards [33].
-Unlearning Scenario We consider two standard settings: i) Single unlearning: A trained model $\mathcal{M}$ receives exactly one request to remove $\mathcal{D}_{f}\subset\mathcal{D}$, producing $\mathcal{M}_{f}=\mathcal{U}(\mathcal{M},\mathcal{D}_{f})$, and ii) Continual unlearning: The model processes a stream of requests ${\mathcal{D}_{f}^{(1)},\dots,\mathcal{D}_{f}^{(k)}}$, updated iteratively by $\mathcal{M}^{(i+1)}=\mathcal{U}(\mathcal{M}^{(t)},\mathcal{D}_{f}^{(i+1)})$ with $\mathcal{M}^{(0)}=\mathcal{M}$ and $\bigcup_{i=1}^{n}\mathcal{D}_{f}^{(t)}=\mathcal{D}_{f}$.
-This mirrors real-world, incremental removal demands while maintaining parity with the single-step budget.
-For the simple tasks, we benchmark all six algorithms: GA, GA + GD, GA + KL, RLabel, NPO, and NPO+KL. For the complex one, lacking a clearly defined retain set, we use GA, NPO, and RLabel.
-Evaluation Metrics For single unlearning (simple tasks only), we report: forget-set accuracy (F.Acc), retain-set accuracy (R.Acc), and privacy leakage via min-$k\%$-prob MIA AUC [27].
+We adopt two open-source LLMs: Yi-6B [41] and Qwen-2.5-7B [39].
+To assess the generality of our findings, we employ two distinct dataset types for unlearning: (i) simple tasks, comprising arXiv abstracts and GitHub code from [40], and (ii) a complex task, NuminaMath-1.5, a recent benchmark for mathematical reasoning [18].
+All experiments are performed on NVIDIA H100 GPUs.
+(Additional results on TOFU [27] and the Traditional-Chinese corpus [^2] https://huggingface.co/datasets/taide/taide-bench are in Appendix.)
+Table 2: Yi-6B: MIA / F.Acc / R.Acc (%) simple task using three LRs under single unlearning
 […]
-Relearning setting.
-To gauge how readily forgotten knowledge re-emerges, each unlearning run is followed by a controlled relearning phase.
-Single unlearning: We fine-tune on the whole forget set $\mathcal{D}_{f}$ once, producing a single-step relearned model.
-Continual unlearning: For settings that trigger catastrophic collapse, we fine-tune with three cases: i) the cumulative forget set $\bigcup_{i}\mathcal{D}_{f}^{(t)}$, ii) the corresponding retain set $\mathcal{D}_{r}^{(t)}$, and iii) an unrelated auxiliary corpus.
-These progressively relax assumptions about access to the forgotten content, revealing the recovery potential in each case.
-Hyperparameter Configuration.
-To comprehensively evaluate the effects of unlearning, we design multiple hyperparameter configurations that vary both the learning rate and the number of unlearning requests.
-For single unlearning we sweep the learning rate over $LR\in\{3,4,5\}\!\times\!10^{-6}$ while fixing the request count to $N=1$.
-For continual unlearning we vary both knobs: on the simple task (Yi-6B) we test $LR\in\{3,5\}\!\times\!10^{-6}\cup\{3\times 10^{-5}\}$ with $N\in\{6\to 100\}$; on the complex task (Qwen-2.5-7B) we use $LR\in\{3,5\}\!\times\!10^{-6}$ and $3\times 10^{-5}$ together with $N\in\{6\to 100\}$.
-All runs adopt the optimizer settings of [1]: AdamW [22] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\varepsilon=10^{-8}$), a cosine schedule with 10% warm-up followed by decay to 10% of peak, weight decay 0.1, and gradient clipping at 1.0.
-Table 1: Yi-6B: MIA / F.Acc / R.Acc (%) simple task using three LRs under single unlearning
-Phase Method LR=3$\times$10-6 LR=4$\times$10-6 LR=5$\times$10-6
 MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc
@@
 RLabel 68.3 78.8 65.6 68.9 76.4 65.3 68.8 78.9 65.2
-3.2 Token-Level Evaluation Results
-We report quantitative results on both single unlearning and continual unlearning settings using Yi-6B and Qwen-2.5-7B across multiple configurations (Tables 13).
-For completeness, detailed experimental results and additional settings are provided in Appendix A.4.
+Unlearning algorithms. We evaluate six canonical unlearning methods, organized into three families.
+1) Gradient-Ascent (GA) family.
+The unified goal is $\mathcal{L}\;=\;\mathcal{L}_{\text{forget}}\!\bigl(\mathcal{D}_{f}\bigr)+\;\lambda\,\mathcal{L}_{\text{retain}}\!\bigl(\mathcal{D}_{r}\bigr),$ where $\mathcal{L}_{\mathrm{forget}}$ maximizes the loss on the forget set via GA, $\mathcal{L}_{\mathrm{retain}}$ (optional) preserves utility on the retain set, and $\lambda>0$ balances the two.
+Choices for $\mathcal{L}_{\mathrm{retain}}$ give three variants: i) GA ($\mathcal{L}_{\mathrm{retain}}=0$), ii) GA+GD (standard cross-entropy on $\mathcal{D}{r}$), and iii) GA+KL (KL divergence to the reference model on $\mathcal{D}{r}$[40].
+2) Negative Preference Optimization (NPO) family.
+GA is replaced by an NPO loss that penalizes agreement with the forget set [43]: $\mathcal{L}\;=\;\mathcal{L}_{\text{NPO}}\!\bigl(\mathcal{D}_{f}\bigr)+\;\lambda\,\mathcal{L}_{\text{retain}}\!\bigl(\mathcal{D}_{r}\bigr),$ Variants mirror those above: NPO ($\mathcal{L}_{\mathrm{retain}}=0$) and NPO+KL (retain-set KL regularization).
+3) Random Label (RLabel).
+To mimic a model that never saw $\mathcal{D}_{f}$, true labels are replaced with random ones: $\mathcal{L}\;=\;\mathcal{L}_{\text{RLabel}}\!\bigl(\mathcal{D}_{f}\bigr),$ inducing near-uniform predictions without GA/negative rewards [40].
+Unlearning Scenarios.
+We consider two standard settings: i) Single unlearning: A trained model $\mathcal{M}$ receives exactly one request to remove $\mathcal{D}_{f}\subset\mathcal{D}$, and ii) Continual unlearning: The model processes a stream of requests $\mathcal{D}_{f}^{(1)},\dots,\mathcal{D}_{f}^{(t)}$}, yielding a sequence of models where $\mathcal{M}^{(t)}=\mathcal{U}(\mathcal{M}^{(t-1)},\mathcal{D}_{f}^{(t)})$.
+For simple tasks, we benchmark all six algorithms. For the complex math reasoning task, where a well-defined retain set is not available, we evaluate the core GA, NPO, and RLabel methods.
+Evaluation Metrics. In single unlearning (simple tasks), we measure forget-set accuracy (F.Acc), retain-set accuracy (R.Acc), and privacy leakage via min-$k\%$-prob MIA AUC [31].
+[…]
+Relearning Setting.
+To assess the reversibility and of unlearning, each run is followed by a controlled relearning phase.
+The unlearned model is briefly fine-tuned on specific data without access to the full pre-training corpus.
+For single unlearning, we fine-tune once on the entire forget set $\mathcal{D}_{f}$.
+For continual unlearning, we evaluate three conditions: (i) the cumulative forget set $\bigcup_{t}\mathcal{D}_{f}^{(t)}$, representing a worst-case adversarial scenario, (ii) the corresponding retain subset $\mathcal{D}_{r}^{(t)}$, as a proxy for the data distribution, and (iii) unrelated out-of-distribution data (general-domain samples explicitly different from $\mathcal{D}_{f}$).
+Each relearning dataset is size-matched to its corresponding unlearning request.
+Hyperparameter Configuration.
+To comprehensively evaluate the effects of unlearning, we design multiple hyperparameter configurations that vary both the learning rate and the number of unlearning requests.
+For single unlearning we sweep the learning rate over $\mathrm{LR}\in\{3,4,5\}\!\times\!10^{-6}$ while fixing the request count to $N=1$.
+For continual unlearning we vary both knobs: on the simple task (Yi‑6B) we test $\mathrm{LR}\in\{3,5\}\!\times\!10^{-6}\cup\{3\times 10^{-5}\}$ with $N\in\{6\to 100\}$; on the complex task (Qwen‑2.5‑7B) we use $\mathrm{LR}\in\{3,5\}\!\times\!10^{-6}$ and $3\times 10^{-5}$ together with $N\in\{6\to 100\}$.
+All runs adopt the optimizer settings of [34]: AdamW [25] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\varepsilon=10^{-8}$), a cosine schedule with 10% warm‑up followed by decay to 10% of peak, weight decay 0.1, and gradient clipping at 1.0.
+3.2 Evaluation Results
+We report quantitative results for single and continual unlearning on Yi-6B and Qwen-2.5-7B under various configurations. Complete results are provided in Appendix Tables 6 and 7.
 Single Unlearning.
-For Yi-6B under single unlearning, Table 1 shows that all six unlearning methods reduce MIA and F.Acc, indicating some level of unlearning.
-The changes on retain set are modest: R.Acc only drops 2–5% for most methods, and MIA decreases by less than 30 points in most cases.
-Importantly, the relearned models often recover original performance—both GA+KL and RLabel restore R.Acc to near 65.0%, and F.Acc rebounds above 77%.
-These results suggest that under single unlearning, most methods achieve seemingly successful forgetting at the token level, but as we show in Section 4.2 to Section 4.2, the underlying representation changes are minimal—indicating the phenomenon of reversible forgetting.
-Continual Unlearning.
-By examining post-relearning recoverability in Table 2 and Table 3, we identify two distinct forms of catastrophic forgetting.
-When the model regains both utility (e.g., F.Acc, R.Acc) and privacy (e.g., MIA AUC) to levels near or exceeding the original after relearning, we classify the behavior as reversible catastrophic forgetting.
-This suggests that the underlying representational structure remains intact, enabling efficient recovery via lightweight retraining.
-Such reversibility is consistently observed in methods like NPO and NPO+KL, particularly under low learning rates or small removal batches.
-Conversely, when relearning fails to restore utility—reflected in persistently low F.Acc and R.Acc despite partial MIA recovery—we categorize it as irreversible catastrophic forgetting.
-This scenario frequently arises with methods like GA and RLabel under aggressive hyperparameters (e.g., $\mathrm{LR}=3\times 10^{-5}$), where damage accumulates across layers and results in irreversible representational collapse.
-Importantly, MIA AUC alone can be misleading.
-Models may exhibit near-complete privacy recovery while remaining functionally impaired.
-Table 2: Yi-6B: MIA / F.Acc / R.Acc (%) for simple task under four unlearning settings
-Phase Method LR=3$\times$$10^{-5}$, N=100 LR=5$\times$$10^{-6}$, N=100 LR=3$\times$$10^{-6}$, N=100 LR=3$\times$$10^{-5}$, N=6
+On Yi-6B, all six methods successfully reduce MIA and F.Acc, indicating a certain degree of forgetting (Table 2).
+The impact on the retain set is modest, with R.Acc dropping by only 2–5%.
+However, relearning often restores original performance; for instance, GA+KL and RLabel recover R.Acc to approximately 65% and F.Acc above 77%.
+These findings suggest that single unlearning achieves superficial forgetting, as the underlying representations remain largely intact (Section 4.2.1).
+This outcome characterizes the reversible, non-catastrophic forgetting regime.
+Table 3: Yi-6B: MIA / F.Acc / R.Acc (%) for simple task under four unlearning settings. Bold numbers indicate improvements over the Original baseline in F.Acc or R.Acc.
+[…]
 MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc
@@
 GA+KL 70.7 1.7 1.6 68.3 81.1 64.8 70.7 81.0 63.2 65.0 66.6 56.2
-NPO 70.0 57.0 45.6 68.0 82.7 65.5 69.9 81.2 65.4 68.4 71.2 59.4
-NPO+KL 67.7 60.7 54.2 69.5 83.8 65.6 69.9 83.8 65.4 69.0 67.6 56.1
+NPO 70.0 57.0 45.6 68.0 82.7 65.5 69.9 81.2 65.5 68.4 71.2 59.4
+NPO+KL 67.7 60.7 54.2 69.5 83.8 65.6 69.9 83.8 65.5 69.0 67.6 56.1
 RLabel 69.5 4.3 2.8 70.4 80.8 65.3 70.0 80.5 65.3 65.2 72.7 61.1
-Table 3: Qwen-2.5-7B: MIA / MATH / GSM8K Accuracy (%) for complex task under four settings
-Phase Method LR=$3\times 10^{-5}$, $N$=6 LR=$3\times 10^{-6}$, $N$=6 LR=$5\times 10^{-6}$, $N$=6 LR=$5\times 10^{-6}$, $N$=100
-MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K
-Original —– 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1
-Unlearn GA 5.9 0.0 0.0 0.9 0.0 0.0 3.8 0.0 0.0 5.5 0.0 0.0
-NPO 95.9 0.0 0.2 97.4 21.5 74.1 67.4 24.1 71.8 94.7 0.0 0.4
-RLabel 35.5 0.0 0.0 69.6 0.0 1.5 11.2 0.0 0.0 2.9 0.0 0.0
-Relearn GA 97.6 0.0 1.1 99.3 5.1 83.2 99.4 9.3 77.8 99.2 0.0 0.0
-NPO 95.8 0.0 0.0 99.4 4.7 82.6 99.4 16.5 75.7 99.2 0.0 0.0
-RLabel 99.5 0.0 0.0 99.3 5.3 83.3 99.3 10.0 77.2 99.6 0.0 0.0
-4 A Unified Representational Analysis
+Continual Unlearning.
+Post-relearning analysis (Tables 3, 6, and 7) reveals two forms of reversible forgetting.
+In reversible, catastrophic forgetting, both utility (e.g., F.Acc, R.Acc) and privacy metrics drop sharply during unlearning but are fully restored after relearning.
+This is observed in GA and RLabel with moderate hyperparameters.
+Besides, reversible, non-catastrophic forgetting entails only a mild, easily recoverable performance degradation, as seen with NPO at $\mathrm{LR}=3\times 10^{-5},\,N=6$.
+Conversely, irreversible, catastrophic forgetting occurs when relearning fails to restore utility, leaving F.Acc and R.Acc low despite partial MIA recovery.
+This pattern is common for GA and RLabel under aggressive hyperparameters (e.g., $\mathrm{LR}=3\times 10^{-5},N=100$), where cumulative updates lead to irreversible representational collapse.
+The MIA AUC metric behaves erratically in this regime: it may fall below 50% during unlearning but misleadingly rebounds to high values after relearning, even after the model’s capabilities have been permanently lost.
+These empirical results on single and continual unlearning are consistent with the theoretical framework in Section 5, which shows that small perturbations to the model weights can trigger disproportionately large drops in accuracy.
+4 Representation-level Evaluation
 4.1 Representational Analysis Tools
-We monitor representational drift with four layer-wise diagnostics: PCA Similarity, PCA Shift, CKA, and the diagonal Fisher Information Matrix (FIM), summarized in Figure 1(b); implementation details (including their precise definitions) are deferred to Appendix A.3.
-PCA Similarity & Shift.
-For each layer $i$, we collect activations $\mathbf{H}_{i}^{\text{orig}}$, $\mathbf{H}_{i}^{\text{unl}}$, and $\mathbf{H}_{i}^{\text{rel}}$ on a probe set $\mathcal{X}$.
-Let $\mathbf{c}_{i,1}^{(*)}$ and $p_{i,1}^{(*)}$ be he first principal direction and its mean projection for state (){$(*)\!\in\!\{\text{orig},\text{unl},\text{rel}\}$.
-The cosine between $\mathbf{c}_{i,1}^{\text{orig}}$ and $\mathbf{c}_{i,1}^{(*)}$ yields PCA similarity; the signed difference $p_{i,1}^{(*)}-p_{i,1}^{\text{orig}}$ gives PCA shift.
-Small angles and shifts indicate stable features; otherwise, catastrophic forgetting [36].
+To analyze representational drift, we employ four hidden state diagnostics, as summarized in Figure 1(b). Their precise definitions and implementation details are deferred to Appendix A.3.
+PCA Similarity, Shift, and Mean Distance.
+For each layer $i$, we collect activation matrices $\mathbf{H}_{i}^{\text{orig}}$, $\mathbf{H}_{i}^{\text{unl}}$, and $\mathbf{H}_{i}^{\text{rel}}$ on a probe set $\mathcal{X}$ for the original, unlearned, and relearned models, respectively.
+Let $\mathbf{c}_{i,1}^{(*)}$ and $p_{i,12}^{(*)}$ be the first principal direction and its mean projection for state $(*)\!\in\!\{\text{orig},\text{unl},\text{rel}\}$.
+PCA Similarity is the cosine between $\mathbf{c}_{i,1}^{\text{orig}}$ and $\mathbf{c}_{i,1}^{(*)}$, while PCA Shift is the signed difference $p_{i,12}^{(*)}$.
+Small values for these metrics indicate stable features, whereas large, unrecovered shifts signify irreversible changes [45].
+We also introduce the mean PCA distance, the average Euclidean distance on $p_{i,12}^{(*)}$ across layers, to provide a single scalar measure of representation drift.
 Centered Kernel Alignment (CKA).
-With centered activation matrices $X_{i}^{\text{orig}}$ and $X_{i}^{(*)}$, we compute ${\rm CKA}(X_{i}^{\text{orig}},X_{i}^{(*)})\!\in\![0,1]$; values $\approx 1$ mean nearly identical subspaces, those $\approx 0$ are orthogonal.
-Fisher information.
-We estimate the diagonal empirical FIM by averaging squared gradients over $\mathcal{X}$.
-Comparing $\mathrm{FIM}^{\text{orig}}$, $\mathrm{FIM}^{\text{unl}}$, and $\mathrm{FIM}^{\text{rel}}$ reveals how unlearning flattens the loss landscape and whether relearning restores parameter importance [14; 11].
-We compute all diagnostics not only on the forget set but also on the retain set and on unrelated data, allowing us to distinguish targeted forgetting from broader representational disruption.
+Given centered activation matrices $X_{i}^{\text{orig}}$ and $Y_{i}^{(*)}$, we compute ${\rm CKA}(X_{i}^{\text{orig}},Y_{i}^{(*)})\!\in\![0,1]$.
+Values $\approx 1$ mean nearly identical subspaces, those $\approx 0$ are orthogonal.
+Fisher information (FIM).
+We estimate the diagonal of the empirical FIM by averaging squared gradients over the probe set $\mathcal{X}$.
+Comparing $\mathrm{FIM}^{\text{orig}}$, $\mathrm{FIM}^{\text{unl}}$, and $\mathrm{FIM}^{\text{rel}}$ reveals how unlearning alters the loss landscape and whether relearning restores parameter importance [15; 10].
+All diagnostics are computed not only on the forget set but also on the retain set and unrelated data to distinguish targeted unlearning from general representational degradation.
+4.2 Representational Results
+(a) Simple (Single Unlearning)
+(b) Simple (Single Unlearning)
+(c) Simple (Single Unlearning)
+(d) Simple (Single Unlearning)
+Figure 2: Single unlearning analysis on Yi-6B with GA under a simple task.
+In reversible non-catastrophic forgetting, PCA Similarity, PCA Shift, CKA, and FIM across layers show only minor changes with slight accuracy drops.
+Input queries are drawn from the forget set.
+4.2.1 Single Unlearning
+Figure 2 demonstrates feature-level changes under single unlearning.
+(a) PCA Similarity remains near 1.0 across all layers, with minor, reversible dips, indicating that dominant activation directions are preserved.
+Slight dips in shallow and final layers are rapidly restored after relearning, suggesting minimal and reversible drift.
+(b) PCA shifts are minimal, and relearned representations closely realign with the original.
+(c) CKA values are nearly 1.0 for all model states, confirming that subspace structures remain intact.
+(d) FIM spectra show only mild, temporary shifts that are fully restored after relearning.
+These results, combined with the task-level evaluation in Section 3.2, demonstrate that single unlearning induces reversible, non-catastrophic forgetting.
+This highlights the limitation of classic (task-level) metrics, which fail to capture the superficial nature of the forgetting.
+4.2.2 Continual Unlearning
+As shown in Figures 3 and 5, PCA Similarity and Shift offer complementary views of representational change: similarity reflects global alignment, and shift is more sensitive to local variations.
+Relying on PCA similarity alone can obscure subtle effects; employing both avoids overlooking fine-grained distinctions, enabling a more comprehensive assessment.
+Higher learning rates or more requests cause sharp drops in similarity and large, unrecovered shifts, which is characteristic of irreversible catastrophic forgetting.
+In contrast, milder hyperparameters lead to high similarity and bounded shifts that are restored after relearning, consistent with reversible, catastrophic forgetting.
+This pattern is consistent across probe sets from forget set, retain set, and unrelated data (Figures 10 and 14).
+Figure 4 integrates CKA (top) and FIM (bottom) analyses.
+CKA reveals that mild unlearning maintains stable alignment that recovers post-relearning, while aggressive unlearning causes irreversible degradation.
+The FIM spectra complement this by showing that continual unlearning flattens the loss landscape.
+Extreme hyperparameters induce a permanent leftward shift in sensitivity distributions, whereas moderate settings permit recovery.
+Together, these diagnostics suggest that observed performance loss is often due to temporary suppression rather than permanent erasure of knowledge.
+For conciseness, we present results on forget-set queries in the main text; retain-set queries, which yield similar conclusions under catastrophic forgetting, are in Appendix A.5 (e.g., Figures 12 and 10).
+Mean PCA Distance Analysis.
+To quantify representation-level drift with a single metric, we use the mean PCA distance.
+To assess the metric’s sensitivity, we compute its mean and standard deviation across four random seeds, as well as shuffling the order of unlearning requests.
+As shown in Table 4, higher learning rates consistently increase the mean PCA distance for both unlearned and relearned models At a low learning rate (e.g., $3\times 10^{-6}$), the mean and variance of the distance remain low after relearning, indicating stable and reproducible recovery.
+In contrast, a high learning rate (e.g., $3\times 10^{-5}$) sharply increases both values, reflecting greater variability and incomplete recovery.
+Combining these empirical findings with the task-level results in Section 3, we further connect to the theoretical analysis in Section 5, whose conclusions show that small weight perturbations induce only limited changes in the feature space, thereby enabling recovery, whereas larger perturbations accumulate into substantial representational drift that underlies irreversibility.
 (a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
@@
 (c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(d) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
+(e) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
 (f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-Figure 2: Layer-wise PCA Similarity for GA on Yi-6B (simple task).
-(a–c) vary LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ at $N=100$; (d–f) vary $N\in\{6,50,100\}$ at LR $=3\times 10^{-5}$.
-Similarity near 1 signals reversible (catastrophic) forgetting; sustained low similarity signals irreversible (catastrophic) forgetting.
+Figure 3: Layer-wise PCA Similarity and Shift for GA on Yi-6B (simple task).
+Vary LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ at $N=100$.
+Sustained low similarity or large shifts signal severe, irreversible catastrophic forgetting, whereas partial similarity or small shifts indicate mild, reversible catastrophic forgetting.
+Input queries are drawn from the forget set.
 (a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
@@
 (c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
-(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-Figure 3: PCA Shift for GA on Yi-6B, simple task.
-(a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$.
-Shift magnitude reflects feature displacement: large, unrecovered shifts indicate severe, irreversible (catastrophic) forgetting, while small or fully recovered shifts indicate mild, reversible (catastrophic) forgetting.
-4.2 Representational Results
-Principal Component Analysis: Similarity and Shift.
-Figures 2 and 3 show that, in continual unlearning, larger learning rates or more removal requests drive PCA Similarity sharply downward and leave large, unrecovered PCA Shifts—clear evidence of irreversible catastrophic forgetting.
-Under gentler hyper-parameters, similarity stays high and shifts remain bounded; relearning then realigns both metrics, signalling reversible catastrophic forgetting.
-Remarkably, the choice of relearning data and analyzed data—forget, retain, or even unrelated—makes little difference: all three restore the feature geometry, implying the knowledge was suppressed rather than erased (Figure 10 and  14).
-Due to space limitations, complete figures, single-unlearning results, and additional methods appear in Appendix A.5 and Appendix A.6.
-Centered Kernel Alignment Analysis.
-Figure 4 tracks layer-wise CKA.
-With gentle unlearning, CKA stays near 1 and relearning restores it completely—typical of reversible catastrophic forgetting.
-Stronger updates or many requests push deep-layer CKA sharply lower, and alignment cannot be fully recovered, signalling irreversible catastrophic forgetting.
-The choice of relearning or probe data—forget, retain, or unrelated—barely affects this outcome: once relearning begins, latent structure resurfaces regardless of input (Figure 18).
-Due to space limitations, complete figures and additional settings appear in Appendix A.5 and Appendix A.6.
-(a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(b) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
-(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+[…]
+[…]
+[…]
 Figure 4: CKA for GA on Yi-6B, simple task.
-(a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$.
-High CKA (near 1) indicates strong subspace alignment and reversible (catastrophic) forgetting, whereas low CKA (near 0) reflects severe representational drift and irreversible (catastrophic) forgetting.
-Fisher Information Analysis.
-Continual unlearning progressively flattens the loss landscape (Figures 5.
-Higher learning rates or larger request counts push the FIM spectra sharply left, and—at extreme settings—the shift persists after relearning, signalling irreversible catastrophic forgetting.
-Under gentler hyper-parameters the spectra recentre, indicating reversible catastrophic forgetting.
-Relearning with forget, retain, or unrelated data realigns the FIM almost equally well, confirming that the lost sensitivity is suppressed rather than erased (Figure 34-38).
-Due to space limitations, complete figures and additional settings appear in Appendix A.5 and Appendix A.6.
-(a) LR=$3\times 10^{-6}$, Layer 31
-(b) LR=$5\times 10^{-6}$, Layer 31
-(c) LR=$3\times 10^{-5}$, Layer 31
-(d) $N=6$, Layer 31
-(e) $N=50$, Layer 31
-(f) $N=100$, Layer 31
-Figure 5: FIM (layer 31) for GA on Yi-6B, simple task.
-(a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$.
-Larger leftward shifts indicate a greater flattening of the loss landscape and irreversible (catastrophic) forgetting; spectra that recenter on the original peak denote reversible (catastrophic) forgetting.
-4.3 Representational Theoretical Analysis
-To explain the empirical distinction between reversible and irreversible (catastrophic) forgetting, we present a perturbation model linking unlearning updates to structural collapse across layers.
-Consider an $L$-layer feedforward neural network $f(x)=\sigma(W_{L}\,\sigma(\cdots\sigma(W_{1}x)\cdots))$, with activations $\sigma$ and weights ${W_{i}}_{i=1}^{L}$.
-Unlearning is modeled as layer-wise perturbations $\widetilde{W}_{i}=W_{i}+E_{i}$ with $|E_{i}|=O(\mathrm{LR},N)$, where $\mathrm{LR}$ is the learning rate and $N$ the number of unlearning steps.
-A Neumann-series expansion yields $\widetilde{f}(x)-f(x)=\sum_{\emptyset\neq S\subseteq\{1,\dots,L\}}(W_{L}\circ\cdots\circ E_{i_{k}}\circ\cdots\circ W_{1})(x)$.
-When small perturbations are confined to a few layers, first-order terms dominate, and the effect is reversible (catastrophic) forgetting.
-In contrast, comparable perturbations spread across many layers, higher-order terms accumulate, producing irreversible (catastrophic) forgetting.
+Vary LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$.
+High CKA ( 1) and concentrated FIM spectra indicates reversible catastrophic forgetting, while persistently low CKA and large-shifted, flattened spectra denote severe representational drift and irreversible catastrophic forgetting.
+Input queries are drawn from the forget set.
+Table 4: Yi-6B (GA): Mean $\pm$ standard deviation of PCA distance on forget and retain sets across varying learning rates and random seeds; the number of unlearning requests is fixed.
+Model Learning Rate Phase Seed Mean PCA distance (forget set) Mean PCA distance (retain set)
+[…]
+$\mathbf{3\times 10^{-6}}$ Relearn 12, 22, 32, 42 $2.11\pm 1.42$ $1.64\pm 1.12$
+$\mathbf{5\times 10^{-6}}$ Unlearn 12, 22, 32, 42 $11.52\pm 6.19$ $8.79\pm 5.20$
+$\mathbf{5\times 10^{-6}}$ Relearn 12, 22, 32, 42 $1.37\pm 0.74$ $1.05\pm 0.58$
+$\mathbf{3\times 10^{-5}}$ Unlearn 12, 22, 32, 42 $133.20\pm 45.81$ $121.45\pm 38.60$
+$\mathbf{3\times 10^{-5}}$ Relearn 12, 22, 32, 42 $104.58\pm 39.70$ $95.34\pm 32.40$
+5 Theoretical Analysis
+5.1 A Perturbation Model of Unlearning
+To explain the empirical distinction between reversible and irreversible (catastrophic) forgetting, we introduce a perturbation model that links unlearning updates to representational changes across layers.
+Consider an $L$-layer feedforward network $f(x)=\sigma(W_{L}\,\sigma(\cdots\sigma(W_{1}x)\cdots))$ with activation $\sigma$ and weights ${W_{i}}_{i=1}^{L}$.
+We model unlearning as a layer-wise perturbation $\widetilde{W}_{i}=W_{i}+E_{i}$, where the magnitude of the error term, $|E_{i}|=O(\mathrm{LR},N)$, scales with the learning rate ($\mathrm{LR}$) and the number of unlearning requests $N$.
+A Neumann-series expansion of the network’s output shows that the total change $\widetilde{f}(x)-f(x)$ is defined as $\sum_{\emptyset\neq S\subseteq\{1,\dots,L\}}(W_{L}\circ\cdots\circ E_{i_{k}}\circ\cdots\circ W_{1})(x)$.
+When perturbations are small and localized to a few layers, first-order terms dominate, leading to reversible (catastrophic) forgetting.
+In contrast, when comparable perturbations are distributed across many layers, higher-order interaction terms accumulate, causing structural degradation that results in irreversible (catastrophic) forgetting.
+We can formalize the impact on our diagnostic tools:
 PCA Similarity.
 Let $X_{i}$ and $Y_{i}=X_{i}+E^{\prime}_{i}$ be the centered activations at layer $i$ before and after unlearning.
-By Davis–Kahan theorem [5], $\cos\angle(\mathbf{c}_{i}^{\text{orig}},\mathbf{c}_{i}^{\text{upd}})\approx 1-O(\|E_{i}^{\prime}\|/(\lambda_{1,i}-\lambda_{2,i})),$ with top two eigenvalues $\lambda_{1,i},\lambda_{2,i}$.
+By the Davis–Kahan theorem [5], $\cos\angle(\mathbf{c}_{i}^{\text{orig}},\mathbf{c}_{i}^{\text{upd}})\approx 1-O(\|E_{i}^{\prime}\|/(\lambda_{1,i}-\lambda_{2,i})),$ with top two eigenvalues $\lambda_{1,i},\lambda_{2,i}$.
 The layer-averaged PCA similarity is $\bar{S}_{\mathrm{PCA}}\approx 1-O((1/L)\sum_{i}\|E^{\prime}_{i}\|)$.
 PCA Shift.
-Along the first principal component, the activation-centroid shift is $\Delta p_{i}=\mu_{i,1}^{\text{upd}}-\mu_{i,1}^{\text{orig}}=O(\|E_{i}^{\prime}\|)$.
-Large perturbations $\|E_{i}^{\prime}\|$ spanning many layers cause irreversible representational drift; otherwise, the shifts remain localized and reversible.
+Along the first principal component, the activation-centroid shift is expressed as $p_{i,12}=O(\|E_{i}^{\prime}\|)$.
+Large perturbations $\|E_{i}^{\prime}\|$ propagating across multiple layers lead to irreversible representational drift, whereas smaller perturbations remain localized and thus reversible.
 CKA.
-Let $\widetilde{K}_{Y_{i}}=\widetilde{K}_{X_{i}}+\Delta K_{i}$ be the perturbed Gram matrix.
-Then, $\mathrm{CKA}_{i}$ is computed as $1-O(\|\Delta K_{i}\|_{*}/\|\widetilde{K}_{X_{i}}\|_{*})$, which implies that $\bar{C}\approx 1-O((1/L)\sum_{i}\|\Delta K_{i}\|_{*})$.
+Let $\widetilde{K}_{Y_{i}}=\widetilde{K}_{X_{i}}+\Delta K_{i}$ denote the perturbed Gram matrix at layer $i$.
+The corresponding CKA score is computed as $\mathrm{CKA}_{i}=1-O\!\left(\|\Delta K_{i}\|_{*}/\|\widetilde{K}_{X_{i}}\|_{*}\right)$.
+Averaging across layers yields $\bar{C}\approx 1-O\!\left(\tfrac{1}{L}\sum_{i}\|\Delta K_{i}\|_{*}\right)$, where $\bar{C}$ denotes the layer-averaged CKA.
 Fisher Information.
 Given update $\delta w_{i}=O(\|E_{i}\|)$, the Fisher diagonal behaves as $F_{ii}(w+\delta w)=F_{ii}(w)+O(\|\delta w_{i}\|)$, so the average Fisher becomes $\bar{F}=(1/P)\sum_{i}F_{ii}=F_{0}-O((1/P)\sum_{i}\|E_{i}\|)$.
-Token-level metrics, such as accuracy, MIA, and AUC, may report total collapse even when the model’s internal geometry is largely preserved.
-In reversible forgetting, a few parameter changes (e.g., in output heads or layer norms) can drastically perturb token probabilities while leaving deeper representations intact.
-For the soft-max output $\log p(y|x;\theta+\delta\theta)\approx\log p(y|x;\theta)+\nabla_{\theta}\log p(y|x;\theta)^{\top}\delta\theta+O(\|\delta\theta\|^{2}).$ a small $\delta\theta$ in high-sensitivity regions (near the logits) can dominate the first-order term, producing large drops in accuracy or anomalous AUC scores despite minimal representational drift.
-Unlearning on the forget set applies a weight update $E_{i}=LR\times\,\nabla_{W_{i}}L(D_{f}),$ which both removes over-fitting to $\mathcal{D}_{f}$ and accentuates its principal feature subspace.
-After relearning, parameters may return close to their originals, yet amplified patterns persist, sometimes yielding better performance on an augmented $\mathcal{D}_{f}$ than the baseline.
-Therefore, unlearning can inadvertently act as a contrastive regularizer, further illustrating the mismatch between surface metrics and feature subspaces.
-Summary.
-Reversible (catastrophic) forgetting occurs when perturbations touch only a few layers; PCA similarity/shift, CKA, and FIM remain near baseline.
-In contrast, irreversible (catastrophic) forgetting emerges from large, distributed updates that collapse the model’s representational structure.
-Unlearning acts as a contrastive perturbation: it removes memorized content yet “reinforces” salient features of the forget set, so subsequent relearning can even outperform the original model on related inputs, revealing its “dual role” in both removal and refinement.
-Token-level metrics (e.g., accuracy, MIA) are overly sensitive to small shifts in high-impact parameters and can misclassify the regime.
-Structural diagnostics, complemented by augmented evaluation, can provide a more reliable assessment of whether forgetting is truly irreversible.
-5 Discussion and Takeaways
-Beyond theoretical justifications in Section 4, we summarize main empirical and analytical insights.
-(1) Single vs. continual unlearning, and the role of GA/RLabel.
-Single unlearning rarely produces irreversible collapse: performance is recoverable and representational drift is slight.
-Continual unlearning, especially with large learning rates, often drives the model into permanent failure: $\sim 100$ sequential requests can push both forget- and retain-set accuracy to near zero.
-GA and RLabel already over-forget in the single scenario and magnify this damage when applied continually.
-Adding retain-set terms, as in GA+KL or NPO(+KL) [33; 35; 31], markedly improves stability.
-(2) Collapse stems from structural drift, not true erasure.
-PCA-Similarity/Shift, CKA, and the FIM consistently expose this breakdown: irreversible collapse coincides with large rotations of principal directions, centroid shifts, and vanishing Fisher mass across many layers.
-When perturbations remain local (small LR, few unlearning requests), these diagnostics stay near baseline—reversible forgetting.
-Token-level metrics alone are unreliable: small updates in high-sensitivity parameters (e.g., logits or layer norms) can tank accuracy or inflate MIA AUC while internal geometry is intact.
-(3) Unlearning can act as implicit augmentation.
-In several continual runs, subsequent relearning on the forget set can often yield higher accuracy than that of the original model.
-This surprising outcome suggests that unlearning is not merely a memory deletion mechanism but may also serve as a form of implicit contrastive regularization.
-As detailed in Section 4, unlearning amplifies the feature subspace associated with the forget set, and relearning on augmented inputs can reinforce semantic structure while promoting robustness.
-This process reorganizes internal representations to better capture generalizable patterns, acting as a form of curriculum learning.
-(4) Diagnostics guide irreversible (benign) forgetting.
-Tools such as PCA Similarity and Shift, CKA, and Fisher Information reveal not just the presence of structural drift, but also where and how it arises across layers.
-This enables targeted control: effective unlearning can be guided toward perturbing only those parameters responsible for the forget set, while preserving the representation structure on the retain set and unrelated data.
-This opens a path to targeted, irreversible forgetting, a permanent and isolated removal of information without collateral collapse, offering actionable insights for building safer unlearning algorithms.
+Table 5: Yi-6B (GA+GD+WAGLE) performance under different relearning settings. F.Acc/R.Acc are forget/retain accuracy, with mean PCA distances on the forget and retain sets.
+Phase F.Acc R.Acc Mean PCA distance (forget set) Mean PCA distance (retain set)
+Original model 78.9 65.5 0 0
+[…]
+Unlearn 37.8 55.9 11.84 6.28
+Relearn 46.9 58.3 9.00 5.91
+[…]
+Unlearn 27.8 51.4 26.02 8.37
+Relearn 31.5 53.5 24.56 8.12
+5.2 Bridging Representational Drift and Task-Level Metrics
+Classic (task-level) metrics can be misleading.
+They are highly sensitive to small weights’ changes, particularly in the final layers, which can cause large shifts in output probabilities without altering the model’s deeper representations.
+For a softmax output, a small perturbation $\delta\theta$ to the model parameters yields a large change in log-probability: $\log p(y|x;\theta+\delta\theta)\approx\log p(y|x;\theta)+\nabla_{\theta}\log p(y|x;\theta)^{\top}\delta\theta+O(\|\delta\theta\|^{2})$.
+A minor update to the logits can dominate this first-order term, causing a sharp drop in accuracy that suggests catastrophic forgetting, even if the underlying geometry is preserved.
+This aligns our theoretical model with the empirical findings in Sections 3 and 4.2.
+When $\mathrm{LR}$ or $N$ is small, changes are confined to first-order effects, feature spaces remain intact, and forgetting is reversible.
+When $\mathrm{LR}$ or $N$ is large, higher-order perturbations accumulate across layers, making recovery impossible and leading to irreversible forgetting.
+Figure 3 illustrates such a transition.
+Interestingly, relearning can sometimes lead to performance that exceeds the original model’s accuracy on the forget set (Table 3).
+This suggests that unlearning can act as a form of contrastive regularization, reinforcing salient features related to the forgotten data, which a brief relearning can then exploit.
+5.3 Probing the Limits of Irreversibility
+In our primary experiments, we did not observe irreversible non-catastrophic forgetting; even a small fraction (e.g., $10\%$) of the forget set was sufficient to restore performance.
+To explore this regime, we conducted extra experiments with more constrained relearning conditions.
+We used the GA+GD+WAGLE method [13], which selectively updates influential parameters, and limited the relearning data to either (i) $50\%$ of the retain set or (ii) an equal-sized, unrelated dataset (Table 5).
+Under these conditions, the method exhibited seemingly irreversible, non-catastrophic forgetting.
+The forget set showed large, unrecoverable PCA distances, while the retain set experienced only modest, partially recoverable degradation.
+This demonstrates that achieving the ideal of targeted, permanent unlearning without collateral damage remains an open challenge.
+Defining precise thresholds to distinguish the forgetting regimes (reversible vs.
+irreversible and catastrophic vs.
+non-catastrophic) is non-trivial, as they depend on the unlearning method, task complexity, and other factors.
+Our representational diagnostics reveal structural drift on each layer, pointing toward avenues for more robust and effective unlearning.
+A promising future direction is the design of new loss functions, such as a forget component penalization term to dismantle target representations and a retain component regularization term to preserve the geometry of desired representations.
 6 Conclusion
-We revisit machine unlearning for LLMs through a systematic study of reversibility. Token-level metrics alone can mislead: a model may appear collapsed yet remain fully recoverable.
-To diagnose this gap, we introduce a feature-space toolkit—PCA Similarity, PCA Shift, CKA, and FIM—that cleanly separates reversible from irreversible catastrophic forgetting.
-Our empirical and theoretical results show that true forgetting arises only when many layers undergo coordinated, large-magnitude perturbations; by contrast, minor updates in high-sensitivity regions (e.g., output logits) can slash accuracy or inflate perplexity while leaving internal representations intact.
-These findings call for evaluation protocols that go beyond token-level scores and for algorithms that actively control representational drift.
-We further observe that unlearning, followed by proper relearning, can refine representations and even boost downstream performance.
-Together, these insights chart a path toward safer, more interpretable unlearning in LLMs.
+This work demonstrates that class (task-level) evaluations of LLM unlearning are insufficient, as performance collapse often masks the reversibility of forgetting.
+Models may appear to have erased data while their internal representations remain intact and easily recoverable.
+Our representation-level toolkit reveals that genuine forgetting requires substantial, coordinated weight perturbations.
+Minor updates often create only a superficial, reversible effect.
+We find that achieving the ideal goal of irreversible, non-catastrophic forgetting remains an open challenge, exposing a fundamental limitation in current methods.
+Our findings call for a shift in evaluation, moving beyond surface-level metrics to protocols that measure true representational change.
+This is essential for developing unlearning algorithms that can provide meaningful and trustworthy guarantees of data removal.
 Appendix A Appendix
@@
 The constrained relearning protocol and selected metrics provide clear insights into representational drift but are not exhaustive and do not offer formal privacy guarantees.
-Extending these analyses to diverse architectures, richer unlearning strategies, and rigorous theoretical bounds is an important direction for future work.
-A.2 Broader impacts
-This work provides new tools and theoretical insights for evaluating and guiding machine unlearning in LLMs, with the potential to improve privacy guarantees and regulatory compliance (e.g., GDPR “right to be forgotten”).
-By distinguishing reversible from irreversible forgetting, our diagnostics can help developers ensure that sensitive or copyrighted material is truly expunged rather than merely hidden.
-At the same time, adversaries could exploit reversible forgetting pathways to mask malicious or biased content and then restore it later, highlighting the need for robust defenses.
-The computational overhead of continual unlearning and repeated diagnostics may increase energy consumption, underscoring the importance of efficient implementations.
-Finally, by exposing the fragility of token-level metrics, our work advocates for more trustworthy evaluation standards that balance privacy, utility, and environmental considerations in real-world deployments.
+In this work, we primarily relied on four diagnostic tools—PCA similarity, PCA shift, CKA, and Fisher information—to capture different aspects of representational drift.
+Other feature-level methods, such as correlation-based approaches (e.g., SVCCA [30]), offer similar perspectives on subspace similarity.
+Incorporating a broader suite of analytic tools is an important direction for future work.
+A.2 Related Work
+Machine Unlearning.
+Machine unlearning has emerged as a critical direction for addressing privacy, safety, and bias in large language models (LLMs) [40; 12; 29; 20; 23; 7; 32; 37; 44; 42; 38; 36; 2].
+It is typically defined as either exact or approximate [2].
+Exact unlearning requires the resulting model to be indistinguishable from one retrained from scratch on the retain set, fully eliminating any statistical trace of the forget set.
+Approximate unlearning relaxes this requirement to distributional or behavioral similarity, demanding only comparable outputs (e.g., perplexity or forget-set accuracy) between unlearned and retrained models [27; 32].
+For modern LLMs, however, exact unlearning is computationally infeasible, as full retraining or partition-based schemes scale poorly [2].
+Consequently, approximate methods dominate practice in LLMs.
+Single-Shot Unlearning.
+Most existing approaches are designed for single deletion events.
+Gradient-based strategies (e.g., GA) enforce forgetting directly but often incur significant utility loss [40].
+Recent advances such as WAGLE augment these methods with weight attribution (e.g., GA+GD+WAGLE), selectively updating the most influential parameters to enhance forgetting efficacy while mitigating utility degradation [13].
+Prompt-based steering avoids parameter updates, reducing cost, but typically achieves only superficial forgetting with vulnerability to reactivation [23].
+Model-editing methods, such as AlphaEdit [21], are lightweight and potentially robust, yet their behavior under sequential or heterogeneous requests remains underexplored.
+Continual Unlearning.
+When unlearning requests arrive sequentially, naive extensions of single-shot methods tend to compound damage, leading to catastrophic forgetting and unstable dynamics [1; 32].
+Each request operates on an already modified model, magnifying utility loss.
+Recent work has attempted to mitigate this through orthogonal updates (e.g., LoRA-based unlearning [11]) and OOD detectors.
+ALKN [36] advances this line by providing a principled framework for continual unlearning, introducing parameter-level interventions and adaptive modules to counteract accumulative decline.
+Evaluations.
+Evaluating unlearning efficacy remains an open challenge.
+Existing studies rely on three main classes of metrics.
+First, classic (task-level) metrics such as accuracy, perplexity [40; 20] are widely used but can be misleading, since performance degradation does not guarantee removal of knowledge.
+Second, memorization probes [17] assess verbatim recall, offering finer granularity but failing to capture semantic or paraphrased knowledge.
+Third, robustness-based evaluations examine vulnerabilities to jailbreaking [46], relearning attacks [24], and even quantization attacks [44].
+For quantization attacks, low-bit compression restores forget-set behavior without direct access to the forget.
+While existing approaches expose weaknesses in reversibility, they often conflate forgetting with class task-level degradation and lack structural insight.
+Our representation-level toolkit closes this gap by jointly diagnosing reversibility and catastrophicity, yielding a more faithful understanding of what is truly forgotten.
+We apply this toolkit to both single unlearning and continual unlearning, the latter of which has not been systematically investigated despite being a more realistic scenario where deletion requests arrive sequentially over a model’s lifecycle.
 A.3 Detailed Analysis Tools
@@
 The PCA Similarity is defined as
-$\operatorname{PCA\text{-}Sim}(i)=\cos\bigl(\mathbf{c}_{i,1}^{\text{orig}},\mathbf{c}_{i,1}^{\text{upd}}\bigr{missing})=\frac{(\mathbf{c}_{i,1}^{\text{orig}})^{\top}\mathbf{c}_{i,1}^{\text{upd}}}{\lVert\mathbf{c}_{i,1}^{\text{orig}}\rVert\,\lVert\mathbf{c}_{i,1}^{\text{upd}}\rVert}\in[-1,1],$
-where values near $1$ indicate stable directional alignment, and values near $-1$ suggest a near-orthogonal shift in dominant directions.
+$\operatorname{PCA\text{-}Sim}(i)=\cos\bigl(\mathbf{c}_{i,1}^{\text{orig}},\mathbf{c}_{i,1}^{\text{upd}}\bigr.)=\frac{(\mathbf{c}_{i,1}^{\text{orig}})^{\top}\mathbf{c}_{i,1}^{\text{upd}}}{\lVert\mathbf{c}_{i,1}^{\text{orig}}\rVert\,\lVert\mathbf{c}_{i,1}^{\text{upd}}\rVert}\in[-1,1],$
+where values near $1$ indicate stable directional alignment, and values near 1-1 suggest a near-orthogonal shift in dominant directions.
 To capture translational drift, we also compute the mean projection of activations along PC1 and PC2:
-$\operatorname{PCA\text{-}Shift}(i)=p_{1,\text{upd}}-p_{1,\text{orig}},\qquad\operatorname{Principle}(i)=p_{2,\text{upd}},$
-where $\operatorname{PCA\text{-}Shift}$ quantifies displacement along PC1 and $\operatorname{Principle}$ captures orthogonal deviation along PC2.
-These metrics reflect how the representation center drifts within the top subspace.
+${PC1\ \Delta}(i)={\mathbf{c}_{i,1}}^{\text{upd}}-{\mathbf{c}_{i,1}}^{\text{orig}},\qquad{PC2}(i)={\mathbf{c}_{i,2}}^{\text{upd}},\qquad p_{i,12}=(PC1\ \Delta(i),PC2(i))$
+where $p_{i,12}$} quantifies displacement along PC1 and captures orthogonal deviation along PC2. These metrics reflect how the representation center drifts within the top subspace.
 Centered Kernel Alignment (CKA).
@@
 $\operatorname{CKA}(X,Y)=\frac{\operatorname{Tr}(\widetilde{K}_{X}\widetilde{K}_{Y})}{\sqrt{\operatorname{Tr}(\widetilde{K}_{X}^{2})}\sqrt{\operatorname{Tr}(\widetilde{K}_{Y}^{2})}}\in[0,1],$
-where values near $1$ indicate highly overlapping subspaces, and values near 00 signal near-orthogonality.
+where values near $1$ indicate highly overlapping subspaces, and values near 0 signal near-orthogonality.
 Fisher Information.
@@
 Together, these tools form a feature-space diagnostic suite: FIM captures global sensitivity, CKA measures subspace preservation, and PCA-based metrics expose fine-grained geometric drift across layers—enabling a robust assessment of representational degradation during unlearning.
-A.4 Complete Experiment
 Table 6: Yi-6B simple-task metrics under four $(\mathrm{LR},N)$ settings.
@@
 Phase Method F.Ppl R.Ppl F.Acc R.Acc CSQA GSM8K MIA
-LR=3×$1003\times 10^{-5},\,N=$ 100
+[…]
 Original 3.8 7.8 78.9 65.5 73.1 39.6 70.9
@@
 Rlable 4056.1 15048.6 4.3 2.8 19.7 0.0 69.5
-LR=5×$1005\times 10^{-6},\,N=$ 100
+[…]
 Unlearn GA $\infty$ $\infty$ 9.1 6.2 19.6 0.0 23.2
@@
 Rlable 3.6 7.7 80.8 65.3 71.8 39.2 70.3
-LR=3×$1003\times 10^{-6},\,N=$ 100
+[…]
 Unlearn GA $\infty$ $\infty$ 16.8 14.4 69.5 12.3 25.2
@@
 Rlable 3.6 7.6 80.5 65.3 72.2 39.2 70.0
-LR=$3\times 10^{-5},\,N=6$
+[…]
 Unlearn GA inf inf 36.3 36.1 69.1 5.8 29.6
@@
 Rlable 6.4 11.4 72.7 61.1 67.5 28.9 65.2
-Table 4 summarizes token-level performance under four unlearning configurations on Yi-6B, varying both learning rates and request counts.
-Most metrics—including forget/retain accuracy, perplexity, CSQA, and GSM8K—follow consistent trends: milder methods (e.g., NPO, NPO+KL) yield moderate degradation and support effective recovery via relearning, while aggressive methods (e.g., GA, RL) under high learning rates lead to severe, often irrecoverable performance drops.
-These results align closely with our theoretical distinction: reversible (catastrophic) forgetting corresponds to localized perturbations and structurally recoverable models, while irreversible (catastrophic) forgetting reflects widespread, irreversible damage.
-MIA, however, behaves differently.
-Even when downstream performance collapses, MIA scores often remain high and recover rapidly.
-This suggests MIA is especially sensitive to shallow changes (e.g., output logits or normalization), rather than deeper representational shifts.
-As a result, MIA often correlates with reversible forgetting—capturing surface-level instability without revealing structural collapse.
-These observations reinforce that while token-level metrics offer partial signals, only internal diagnostics can reliably differentiate reversible from irreversible forgetting.
-A.5 Single Unlearning
-Figure 6 illustrates feature-level changes under single unlearning using PCA Similarity, PCA Shift, CKA, and Fisher Information.
-In (a), PCA Similarity remains consistently high across layers, with cosine scores near 1, indicating that dominant activation directions are well preserved.
-Slight dips in shallow and final layers are quickly restored after relearning, suggesting minimal and reversible geometric drift.
-Subfigure (b) confirms that PC1 shifts and orthogonal deviations are small, with relearned centers closely matching the original.
-In (c), CKA shows near-perfect alignment between the original, unlearned, and relearned representations, reinforcing the conclusion that subspace structure remains intact.
-Fisher Information spectra in (d–f) reveal only mild leftward shifts, indicating slight loss flattening and reduced parameter sensitivity, which are fully recovered after relearning.
-Overall, these results confirm that single unlearning causes only minor, reversible structural perturbations—highlighting the fragility of token-level evaluations in capturing irreversible forgetting.
-(a) Simple (Single Unlearning)
-(b) Simple (Single Unlearning)
-(c) Simple (Single Unlearning)
-(d) Simple (Single Unlearning)
-(e) Simple (Single Unlearning)
-(f) Simple (Single Unlearning)
-Figure 6: Single unlearning analysis on Yi-6B for GA under a simple task.
-PCA Similarity, PCA Shift, CKA, and Fisher information (FIM) distributions are reported across key layers to assess representation and parameter changes before and after unlearning and relearning.
-A.6 Detailed Analysis Results
+Table 7: Qwen-2.5-7B: MIA / MATH / GSM8K Accuracy (%) for complex task under four settings. Bold numbers indicate improvements over the Original baseline in MATH or GSM8K.
+[…]
+MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K
+Original —– 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1
+Unlearn GA 5.9 0.0 0.0 0.9 0.0 0.0 3.8 0.0 0.0 5.5 0.0 0.0
+NPO 95.9 0.0 0.2 97.4 21.5 74.1 67.4 24.1 71.8 94.7 0.0 0.4
+RLabel 35.5 0.0 0.0 69.6 0.0 1.5 11.2 0.0 0.0 2.9 0.0 0.0
+Relearn GA 97.6 0.0 1.1 99.3 5.1 83.2 99.4 9.3 77.8 99.2 0.0 0.0
+NPO 95.8 0.0 0.0 99.4 4.7 82.6 99.4 16.5 75.7 99.2 0.0 0.0
+RLabel 99.5 0.0 0.0 99.3 5.3 83.3 99.3 10.0 77.2 99.6 0.0 0.0
+Table 8: Yi-6B (GA): Mean PCA distance under different learning rates. The left block uses China Taiwan for relearning only, while the right block uses TOFU for both unlearning and relearning.
+Relearning with China Taiwan Unlearning + Relearning with TOFU
+Learning Rate Phase Mean PCA distance (forget set) Phase Mean PCA distance (forget set)
+$3\times 10^{-6}$ Unlearn 17.12 Unlearn 0.51
+$3\times 10^{-6}$ Relearn 4.98 Relearn 0.27
+$5\times 10^{-6}$ Unlearn 20.27 Unlearn 2.41
+$5\times 10^{-6}$ Relearn 10.77 Relearn 1.08
+$3\times 10^{-5}$ Unlearn 193.13 Unlearn 11.96
+$3\times 10^{-5}$ Relearn 167.32 Relearn 11.02
+A.4 Mean PCA distance under Different Dataset
+To examine the role of distributional alignment, we evaluate unlearning and relearning under two dataset settings.
+First, we use the TOFU benchmark [27], where both unlearning and relearning occur within the same distribution.
+Second, treating different languages as out-of-distribution (OOD), we include a Traditional-Chinese corpus for relearning.
+This setup enables us to probe whether cross-lingual signals can drive effective recovery, and how their efficacy compares with in-distribution.
+Table 8 confirms that cross-lingual relearning improves the model but achieves less complete restoration than English data: mean PCA distance and related summary metrics move closer to baseline values, yet remain substantially higher.
+Greater linguistic or domain dissimilarity therefore reduces the efficacy of recovery, though partial restoration is still attainable.
+For the TOFU dataset, the overall pattern holds: learning rate and the number of unlearning requests ($N$) effectively regulate feature drift and reversibility.
+However, the representational shifts induced by TOFU are milder than those observed in our simple and complex tasks.
+We attribute this to the smaller and less diverse nature of TOFU’s corpus; many entries are short and contain only author metadata, making its impact on the model’s feature space comparatively limited.
+A.5 Detailed Analysis Results
+A.5.1 Principal Component Analysis: Similarity and Shift
+Across the same hyper-parameter grid, Figure 7 (PCA–Similarity) and Figure 11 (PCA–Shift) provide complementary views of representational drift.
+For GA, higher learning rates drive unlearned states (orange) far from the original (blue), while relearning (green) fails to return, producing long rays of irreversible drift.
+GA+GD narrows the spread but still collapses at $3{\times}10^{-5}$.
+On Qwen-2.5-7B, GA shifts span thousands of PC1 units and drive PC2 to extreme negatives (Figure 13c,f,i), consistent with the multi-layer perturbations predicted in Section 4.
+In complex tasks such as mathematical reasoning, even small perturbations in hidden states can lead to substantial performance differences.
+This is reflected in our PCA–Similarity analysis, where seemingly minor changes in hidden state geometry correspond to meaningful behavioral variations.
+Besides, PCA-Similarity captures global alignment, whereas PCA–Shift highlights fine-grained translational drift.
+This distinction also explains why Figure 9h,i show only moderate misalignment under similarity but reveal pronounced displacements under shift (cf.
+Figure 13).
+Using both metrics thus provides a more complete characterization of reversibility.
+Overall, these results confirm that GA, with or without GD or KL, induces large and often irreversible displacements, whereas NPO variants, and to a lesser extent RLabel, constrain less shifts, consistent with our utility findings.
+A.5.2 Centered Kernel Alignment Analysis
+Figures 1517 report layer-wise linear CKA between the original model and its unlearned or relearned counterparts.
+Across both Yi-6B and Qwen-2.5-7B, GA stands out: as the learning rate or $N$ increases, its CKA curve drops close to zero in most layers and fails to recover, revealing a deep subspace fracture consistent with the irreversible PCA trends.
+GA+GD and GA+KL mitigate this decline to some extent but do not restore full alignment after relearning.
+Task complexity does not alter the ordering but amplifies the differences.
+On the math-heavy Qwen benchmark, GA’s tail layers fall almost to zero at high learning rates, whereas NPO maintains significantly higher alignment.
+Taken together with the PCA-Shift results, these findings show that GA-style objectives consistently break subspace alignment, NPO variants preserve much greater stability, and RLabel induces moderate but partly recoverable distortions.
+A.5.3 Fisher Information Analysis
+Figures 1933 plot the empirical Fisher spectra layer by layer.
+Across both Yi-6B (simple) and Qwen-2.5-7B (complex), GA and its variants exhibit a pronounced leftward shift of the diagonal histogram as LR or $N$ increase.
+The peaks move several orders of magnitude in middle and deep layers, reflecting a flattened loss surface and diminished parameter salience.
+Crucially, these shifts persist after relearning, marking the onset of irreversible forgetting.
+NPO, NPO+KL, and RL produce smaller leftward displacements under moderate LR or $N$, and their Fisher spectra recenter after relearning, indicating primarily reversible drift.
+Under extreme settings (e.g., ${\rm LR}=3\times 10^{-5}$ or $N=100$), these methods also show persistent displacement in some layers, suggesting milder but still irreversible forgetting.
+Figures 14, 10, 18, and 34 examine relearning dynamics when the fine-tuning data and input query are drawn from the forget set, the retain set, or an unrelated data: i) across all sources, the overall trends are similar: alignment can be partially restored, but recovery is consistently weaker with unrelated data, underscoring that effective relearning depends on both the size and the relevance of the training set; ii) the observed behavior also varies with the choice of input queries.
+In the case of reversible catastrophic forgetting, all forget set, retain set, and unrelated data undergo the similar feature drifts.
+(a) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(b) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+Figure 5: Layer-wise PCA Similarity and Shift for GA on Yi-6B (simple task).
+vary $N\in\{6,50,100\}$ at LR $3\times 10^{-5}$.
+Sustained low similarity or large shifts signal severe, irreversible catastrophic forgetting, whereas partial similarity or small shifts indicate mild, reversible catastrophic forgetting.
+Input queries are drawn from the forget set.
+(a) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(b) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(d) $N=6$, Layer 31
+(e) $N=50$, Layer 31
+(f) $N=100$, Layer 31
+Figure 6: CKA and FIM for GA on Yi-6B, simple task.
+Vary LR $3\times 10^{-5}$ with $N\in\{6,50,100\}$.
+High CKA ( 1) and concentrated FIM spectra indicates reversible catastrophic forgetting, while persistently low CKA and large-shifted, flattened spectra denote severe representational drift and irreversible catastrophic forgetting.
+Input queries are drawn from the forget set.
 (a) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
@@
 (o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-[…]
+Figure 7: PCA Similarity Across Layers.
+Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o).
+All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
 (a) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
@@
 (i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-[…]
+Figure 9: PCA Similarity Across Layers.
+Each row shows results under different unlearning methods: GA (a-c) NPO (d–f), Rlable (g–j).
+All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
 (a) Simple (Relearned by forget set)
@@
 (o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-[…]
+Figure 11: PCA Shift Across Layers.
+Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o).
+All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
 (a) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
@@
 Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
-(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-[…]
+(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
+(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
+(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
+(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
+(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
+(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
+(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+Figure 13: PCA Shift Across Layers.
+Each row shows results under different unlearning methods: GA (a-c) NPO (d–f), Rlable (g–j).
+All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
 (a) Simple (Relearned by forget set)
@@
 (d–f): PCA shift is measured using the forget set, retain set, or unrelated data as evaluation input.
-(a) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(b) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(c) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(d) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(e) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(f) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(g) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(h) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(i) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(j) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(k) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(l) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(m) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
-(n) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
-(o) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-[…]
-(a) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(b) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
-(c) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(d) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(e) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
-(f) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(g) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(h) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
-(i) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(j) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(k) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
-(l) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
-(m) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(n) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
-(o) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(a) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
+(b) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
+(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(d) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
+(e) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
+(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(g) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
+(h) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
+(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(j) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
+(k) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
+(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(m) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
+(n) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
+(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+Figure 15: CKA Across Layers.
+Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o).
+All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
+(a) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(b) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(d) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(e) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(g) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(h) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(j) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(k) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
+(m) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(n) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
+(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
 Figure 16: CKA Across Layers.
@@
 Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
-(a) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
-(b) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
-(c) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(d) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
-(e) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
-(f) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-(g) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
-(h) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
-(i) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
-[…]
+(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
+(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
+(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
+(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
+(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
+(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
+(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
+Figure 17: CKA Across Layers.
+Each row shows results under different unlearning methods: GA (a-c) NPO (d–f), Rlable (g–j).
+All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
 (a) Simple (Relearned by forget set)
@@
 (d–f): CKA is measured using the forget set, retain set, or unrelated data as evaluation input.
-(a) LR=$3\times 10^{-6}$, Layer 28
-(b) LR=$5\times 10^{-6}$, Layer 28
-(c) LR=$3\times 10^{-5}$, Layer 28
-(d) LR=$3\times 10^{-6}$, Layer 22
-(e) LR=$5\times 10^{-6}$, Layer 22
-(f) LR=$3\times 10^{-5}$, Layer 22
-(g) LR=$3\times 10^{-6}$, Layer 13
-(h) LR=$5\times 10^{-6}$, Layer 13
-(i) LR=$3\times 10^{-5}$, Layer 13
-(j) LR=$3\times 10^{-6}$, Layer 4
-(k) LR=$5\times 10^{-6}$, Layer 4
-(l) LR=$3\times 10^{-5}$, Layer 4
-(m) LR=$3\times 10^{-6}$, Layer 1
-(n) LR=$5\times 10^{-6}$, Layer 1
-(o) LR=$3\times 10^{-5}$, Layer 1
 […]
-(a) LR=$3\times 10^{-6}$, Layer 31
-(b) LR=$5\times 10^{-6}$, Layer 31
-(c) LR=$3\times 10^{-5}$, Layer 31
-(d) LR=$3\times 10^{-6}$, Layer 28
-(e) LR=$5\times 10^{-6}$, Layer 28
-(f) LR=$3\times 10^{-5}$, Layer 28
-(g) LR=$3\times 10^{-6}$, Layer 22
-(h) LR=$5\times 10^{-6}$, Layer 22
-(i) LR=$3\times 10^{-5}$, Layer 22
-(j) LR=$3\times 10^{-6}$, Layer 13
-(k) LR=$5\times 10^{-6}$, Layer 13
-(l) LR=$3\times 10^{-5}$, Layer 13
-(m) LR=$3\times 10^{-6}$, Layer 4
-(n) LR=$5\times 10^{-6}$, Layer 4
-(o) LR=$3\times 10^{-5}$, Layer 4
-(p) LR=$3\times 10^{-6}$, Layer 1
-(q) LR=$5\times 10^{-6}$, Layer 1
-(r) LR=$3\times 10^{-5}$, Layer 1
 […]
-(a) LR=$3\times 10^{-6}$, Layer 31
-(b) LR=$5\times 10^{-6}$, Layer 31
-(c) LR=$3\times 10^{-5}$, Layer 31
-(d) LR=$3\times 10^{-6}$, Layer 28
-(e) LR=$5\times 10^{-6}$, Layer 28
-(f) LR=$3\times 10^{-5}$, Layer 28
-(g) LR=$3\times 10^{-6}$, Layer 22
-(h) LR=$5\times 10^{-6}$, Layer 22
-(i) LR=$3\times 10^{-5}$, Layer 22
-(j) LR=$3\times 10^{-6}$, Layer 13
-(k) LR=$5\times 10^{-6}$, Layer 13
-(l) LR=$3\times 10^{-5}$, Layer 13
-(m) LR=$3\times 10^{-6}$, Layer 4
-(n) LR=$5\times 10^{-6}$, Layer 4
-(o) LR=$3\times 10^{-5}$, Layer 4
-(p) LR=$3\times 10^{-6}$, Layer 1
-(q) LR=$5\times 10^{-6}$, Layer 1
-(r) LR=$3\times 10^{-5}$, Layer 1
 […]
-(a) LR=$3\times 10^{-6}$, Layer 31
-(b) LR=$5\times 10^{-6}$, Layer 31
-(c) LR=$3\times 10^{-5}$, Layer 31
-(d) LR=$3\times 10^{-6}$, Layer 28
-(e) LR=$5\times 10^{-6}$, Layer 28
-(f) LR=$3\times 10^{-5}$, Layer 28
-(g) LR=$3\times 10^{-6}$, Layer 22
-(h) LR=$5\times 10^{-6}$, Layer 22
-(i) LR=$3\times 10^{-5}$, Layer 22
-(j) LR=$3\times 10^{-6}$, Layer 13
-(k) LR=$5\times 10^{-6}$, Layer 13
-(l) LR=$3\times 10^{-5}$, Layer 13
-(m) LR=$3\times 10^{-6}$, Layer 4
-(n) LR=$5\times 10^{-6}$, Layer 4
-(o) LR=$3\times 10^{-5}$, Layer 4
-(p) LR=$3\times 10^{-6}$, Layer 1
-(q) LR=$5\times 10^{-6}$, Layer 1
-(r) LR=$3\times 10^{-5}$, Layer 1
 […]
-(a) LR=$3\times 10^{-6}$, Layer 31
-(b) LR=$5\times 10^{-6}$, Layer 31
-(c) LR=$3\times 10^{-5}$, Layer 31
-(d) LR=$3\times 10^{-6}$, Layer 28
-(e) LR=$5\times 10^{-6}$, Layer 28
-(f) LR=$3\times 10^{-5}$, Layer 28
-(g) LR=$3\times 10^{-6}$, Layer 22
-(h) LR=$5\times 10^{-6}$, Layer 22
-(i) LR=$3\times 10^{-5}$, Layer 22
-(j) LR=$3\times 10^{-6}$, Layer 13
-(k) LR=$5\times 10^{-6}$, Layer 13
-(l) LR=$3\times 10^{-5}$, Layer 13
-(m) LR=$3\times 10^{-6}$, Layer 4
-(n) LR=$5\times 10^{-6}$, Layer 4
-(o) LR=$3\times 10^{-5}$, Layer 4
-(p) LR=$3\times 10^{-6}$, Layer 1
-(q) LR=$5\times 10^{-6}$, Layer 1
-(r) LR=$3\times 10^{-5}$, Layer 1
 […]
-(a) LR=$3\times 10^{-6}$, Layer 31
-(b) LR=$5\times 10^{-6}$, Layer 31
-(c) LR=$3\times 10^{-5}$, Layer 31
-(d) LR=$3\times 10^{-6}$, Layer 28
-(e) LR=$5\times 10^{-6}$, Layer 28
-(f) LR=$3\times 10^{-5}$, Layer 28
-(g) LR=$3\times 10^{-6}$, Layer 22
-(h) LR=$5\times 10^{-6}$, Layer 22
-(i) LR=$3\times 10^{-5}$, Layer 22
-(j) LR=$3\times 10^{-6}$, Layer 13
-(k) LR=$5\times 10^{-6}$, Layer 13
-(l) LR=$3\times 10^{-5}$, Layer 13
-(m) LR=$3\times 10^{-6}$, Layer 4
-(n) LR=$5\times 10^{-6}$, Layer 4
-(o) LR=$3\times 10^{-5}$, Layer 4
-(p) LR=$3\times 10^{-6}$, Layer 1
-(q) LR=$5\times 10^{-6}$, Layer 1
-(r) LR=$3\times 10^{-5}$, Layer 1
 […]
-(a) $N=6$, Layer 28
-(b) $N=50$, Layer 28
-(c) $N=100$, Layer 28
-(d) $N=6$, Layer 22
-(e) $N=50$, Layer 22
-(f) $N=100$, Layer 22
-(g) $N=6$, Layer 13
-(h) $N=50$, Layer 13
-(i) $N=100$, Layer 13
-(j) $N=6$, Layer 4
-(k) $N=50$, Layer 4
-(l) $N=100$, Layer 4
-(m) $N=6$, Layer 1
-(n) $N=50$, Layer 1
-(o) $N=100$, Layer 1
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 19: FIM for GA Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 20: FIM for GA+GD Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 21: FIM for GA+KL Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 22: FIM for NPO Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 23: FIM for NPO+KL Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 24: FIM for Rlable Across Layers. All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
+(a) Simple $N=6$, Layer 28
+(b) Simple $N=50$, Layer 28
+(c) Simple $N=100$, Layer 28
+(d) Simple $N=6$, Layer 22
+(e) Simple $N=50$, Layer 22
+(f) Simple $N=100$, Layer 22
+(g) Simple $N=6$, Layer 13
+(h) Simple $N=50$, Layer 13
+(i) Simple $N=100$, Layer 13
+(j) Simple $N=6$, Layer 4
+(k) Simple $N=50$, Layer 4
+(l) Simple $N=100$, Layer 4
+(m) Simple $N=6$, Layer 1
+(n) Simple $N=50$, Layer 1
+(o) Simple $N=100$, Layer 1
 Figure 25: FIM for GA Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
-(a) $N=6$, Layer 31
-(b) $N=50$, Layer 31
-(c) $N=100$, Layer 31
-(d) $N=6$, Layer 28
-(e) $N=50$, Layer 28
-(f) $N=100$, Layer 28
-(g) $N=6$, Layer 22
-(h) $N=50$, Layer 22
-(i) $N=100$, Layer 22
-(j) $N=6$, Layer 13
-(k) $N=50$, Layer 13
-(l) $N=100$, Layer 13
-(m) $N=6$, Layer 4
-(n) $N=50$, Layer 4
-(o) $N=100$, Layer 4
-(p) $N=6$, Layer 1
-(q) $N=50$, Layer 1
-(r) $N=100$, Layer 1
+(a) Simple $N=6$, Layer 31
+(b) Simple $N=50$, Layer 31
+(c) Simple $N=100$, Layer 31
+(d) Simple $N=6$, Layer 28
+(e) Simple $N=50$, Layer 28
+(f) Simple $N=100$, Layer 28
+(g) Simple $N=6$, Layer 22
+(h) Simple $N=50$, Layer 22
+(i) Simple $N=100$, Layer 22
+(j) Simple $N=6$, Layer 13
+(k) Simple $N=50$, Layer 13
+(l) Simple $N=100$, Layer 13
+(m) Simple $N=6$, Layer 4
+(n) Simple $N=50$, Layer 4
+(o) Simple $N=100$, Layer 4
+(p) Simple $N=6$, Layer 1
+(q) Simple $N=50$, Layer 1
+(r) Simple $N=100$, Layer 1
 Figure 26: FIM for GA+GD Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
-(a) $N=6$, Layer 31
-(b) $N=50$, Layer 31
-(c) $N=100$, Layer 31
-(d) $N=6$, Layer 28
-(e) $N=50$, Layer 28
-(f) $N=100$, Layer 28
-(g) $N=6$, Layer 22
-(h) $N=50$, Layer 22
-(i) $N=100$, Layer 22
-(j) $N=6$, Layer 13
-(k) $N=50$, Layer 13
-(l) $N=100$, Layer 13
-(m) $N=6$, Layer 4
-(n) $N=50$, Layer 4
-(o) $N=100$, Layer 4
-(p) $N=6$, Layer 1
-(q) $N=50$, Layer 1
-(r) $N=100$, Layer 1
+(a) Simple $N=6$, Layer 31
+(b) Simple $N=50$, Layer 31
+(c) Simple $N=100$, Layer 31
+(d) Simple $N=6$, Layer 28
+(e) Simple $N=50$, Layer 28
+(f) Simple $N=100$, Layer 28
+(g) Simple $N=6$, Layer 22
+(h) Simple $N=50$, Layer 22
+(i) Simple $N=100$, Layer 22
+(j) SimpleN=6N=6, Layer 13
+(k) Simple $N=50$, Layer 13
+(l) Simple $N=100$, Layer 13
+(m) Simple $N=6$, Layer 4
+(n) Simple $N=50$, Layer 4
+(o) Simple $N=100$, Layer 4
+(p) Simple $N=6$, Layer 1
+(q) Simple $N=50$, Layer 1
+(r) Simple $N=100$, Layer 1
 Figure 27: FIM for GA+KL Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
-(a) $N=6$, Layer 31
-(b) $N=50$, Layer 31
-(c) $N=100$, Layer 31
-(d) $N=6$, Layer 28
-(e) $N=50$, Layer 28
-(f) $N=100$, Layer 28
-(g) $N=6$, Layer 22
-(h) $N=50$, Layer 22
-(i) $N=100$, Layer 22
-(j) $N=6$, Layer 13
-(k) $N=50$, Layer 13
-(l) $N=100$, Layer 13
-(m) $N=6$, Layer 4
-(n) $N=50$, Layer 4
-(o) $N=100$, Layer 4
-(p) $N=6$, Layer 1
-(q) $N=50$, Layer 1
-(r) $N=100$, Layer 1
+(a) Simple $N=6$, Layer 31
+(b) Simple $N=50$, Layer 31
+(c) Simple $N=100$, Layer 31
+(d) SimpleN=6N=6, Layer 28
+(e) Simple $N=50$, Layer 28
+(f) Simple $N=100$, Layer 28
+(g) Simple $N=6$, Layer 22
+(h) Simple $N=50$, Layer 22
+(i) Simple $N=100$, Layer 22
+(j) Simple $N=6$, Layer 13
+(k) Simple $N=50$, Layer 13
+(l) Simple $N=100$, Layer 13
+(m) Simple $N=6$, Layer 4
+(n) Simple $N=50$, Layer 4
+(o) Simple $N=100$, Layer 4
+(p) Simple $N=6$, Layer 1
+(q) Simple $N=50$, Layer 1
+(r) Simple $N=100$, Layer 1
 Figure 28: FIM for NPO Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
-(a) $N=6$, Layer 31
-(b) $N=50$, Layer 31
-(c) $N=100$, Layer 31
-(d) $N=6$, Layer 28
-(e) $N=50$, Layer 28
-(f) $N=100$, Layer 28
-(g) $N=6$, Layer 22
-(h) $N=50$, Layer 22
-(i) $N=100$, Layer 22
-(j) $N=6$, Layer 13
-(k) $N=50$, Layer 13
-(l) $N=100$, Layer 13
-(m) $N=6$, Layer 4
-(n) $N=50$, Layer 4
-(o) $N=100$, Layer 4
-(p) $N=6$, Layer 1
-(q) $N=50$, Layer 1
-(r) $N=100$, Layer 1
+(a) Simple $N=6$, Layer 31
+(b) Simple $N=50$, Layer 31
+(c) Simple $N=100$, Layer 31
+(d) Simple $N=6$, Layer 28
+(e) Simple $N=50$, Layer 28
+(f) Simple $N=100$, Layer 28
+(g) Simple $N=6$, Layer 22
+(h) Simple $N=50$, Layer 22
+(i) Simple $N=100$, Layer 22
+(j) Simple $N=6$, Layer 13
+(k) Simple $N=50$, Layer 13
+(l) SimpleN=100N=100, Layer 13
+(m) Simple $N=6$, Layer 4
+(n) Simple $N=50$, Layer 4
+(o) Simple $N=100$, Layer 4
+(p) Simple $N=6$, Layer 1
+(q) Simple $N=50$, Layer 1
+(r) SimpleN=100N=100, Layer 1
 Figure 29: FIM for NPO+KL Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
-(a) $N=6$, Layer 31
-(b) $N=50$, Layer 31
-(c) $N=100$, Layer 31
-(d) $N=6$, Layer 28
-(e) $N=50$, Layer 28
-(f) $N=100$, Layer 28
-(g) $N=6$, Layer 22
-(h) $N=50$, Layer 22
-(i) $N=100$, Layer 22
-(j) $N=6$, Layer 13
-(k) $N=50$, Layer 13
-(l) $N=100$, Layer 13
-(m) $N=6$, Layer 4
-(n) $N=50$, Layer 4
-(o) $N=100$, Layer 4
-(p) $N=6$, Layer 1
-(q) $N=50$, Layer 1
-(r) $N=100$, Layer 1
+(a) Simple $N=6$, Layer 31
+(b) Simple $N=50$, Layer 31
+(c) Simple $N=100$, Layer 31
+(d) Simple $N=6$, Layer 28
+(e) Simple $N=50$, Layer 28
+(f) Simple $N=100$, Layer 28
+(g) Simple $N=6$, Layer 22
+(h) Simple $N=50$, Layer 22
+(i) Simple $N=100$, Layer 22
+(j) Simple $N=6$, Layer 13
+(k) Simple $N=50$, Layer 13
+(l) Simple $N=100$, Layer 13
+(m) Simple $N=6$, Layer 4
+(n) Simple $N=50$, Layer 4
+(o) Simple $N=100$, Layer 4
+(p) Simple $N=6$, Layer 1
+(q) Simple $N=50$, Layer 1
+(r) Simple $N=100$, Layer 1
 Figure 30: FIM for Rlable Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
-(a) LR=$3\times 10^{-6}$, Layer 28
-(b) LR=$5\times 10^{-6}$, Layer 28
-(c) LR=$3\times 10^{-5}$, Layer 28
-(d) LR=$3\times 10^{-6}$, Layer 24
-(e) LR=$5\times 10^{-6}$, Layer 24
-(f) LR=$3\times 10^{-5}$, Layer 24
-(g) LR=$3\times 10^{-6}$, Layer 12
-(h) LR=$5\times 10^{-6}$, Layer 12
-(i) LR=$3\times 10^{-5}$, Layer 12
-(j) LR=$3\times 10^{-6}$, Layer 4
-(k) LR=$5\times 10^{-6}$, Layer 4
-(l) LR=$3\times 10^{-5}$, Layer 4
-(m) LR=$3\times 10^{-6}$, Layer 1
-(n) LR=$5\times 10^{-6}$, Layer 1
-(o) LR=$3\times 10^{-5}$, Layer 1
 […]
-(a) LR=$3\times 10^{-6}$, Layer 28
-(b) LR=$5\times 10^{-6}$, Layer 28
-(c) LR=$3\times 10^{-5}$, Layer 28
-(d) LR=$3\times 10^{-6}$, Layer 24
-(e) LR=$5\times 10^{-6}$, Layer 24
-(f) LR=$3\times 10^{-5}$, Layer 24
-(g) LR=$3\times 10^{-6}$, Layer 12
-(h) LR=$5\times 10^{-6}$, Layer 12
-(i) LR=$3\times 10^{-5}$, Layer 12
-(j) LR=$3\times 10^{-6}$, Layer 4
-(k) LR=$5\times 10^{-6}$, Layer 4
-(l) LR=$3\times 10^{-5}$, Layer 4
-(m) LR=$3\times 10^{-6}$, Layer 1
-(n) LR=$5\times 10^{-6}$, Layer 1
-(o) LR=$3\times 10^{-5}$, Layer 1
 […]
-(a) LR=$3\times 10^{-6}$, Layer 28
-(b) LR=$5\times 10^{-6}$, Layer 28
-(c) LR=$3\times 10^{-5}$, Layer 28
-(d) LR=$3\times 10^{-6}$, Layer 24
-(e) LR=$5\times 10^{-6}$, Layer 24
-(f) LR=$3\times 10^{-5}$, Layer 24
-(g) LR=$3\times 10^{-6}$, Layer 12
-(h) LR=$5\times 10^{-6}$, Layer 12
-(i) LR=$3\times 10^{-5}$, Layer 12
-(j) LR=$3\times 10^{-6}$, Layer 4
-(k) LR=$5\times 10^{-6}$, Layer 4
-(l) LR=$3\times 10^{-5}$, Layer 4
-(m) LR=$3\times 10^{-6}$, Layer 1
-(n) LR=$5\times 10^{-6}$, Layer 1
-(o) LR=$3\times 10^{-5}$, Layer 1
 […]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 31: FIM for GA Across Layers. All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 32: FIM for NPO Across Layers. All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 33: FIM for Rlable Across Layers. All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
 (a) Simple (Relearned by forget set), Layer 31
@@
 (c) Simple (Relearned by unrelated data), Layer 4
-(d) Simple (input data = forget set Layer), 4
-(e) Simple (input data = retain set Layer), 4
+(d) Simple (input data = forget set), Layer 4
+(e) Simple (input data = retain set), Layer 4
 (f) Simple (input data = unrelated data), Layer 4
@@
 (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
-A.6.1 Principal Component Analysis: Similarity and Shift
-Across the same hyper-parameter grid, Figures 7– 13 plot the PCA–Shift trajectories—layer-wise displacements of activation centroids along the first two principal directions.
-For GA-based objectives the pattern mirrors their Similarity curves: as LR rises the orange triangles (unlearned) shoot far from the blue circles (original), especially in deeper layers, and the green squares (relearned) fail to return, producing long grey rays that diagnose irreversible drift.
-With GA+GD the spread is smaller but still widens sharply at $3{\times}10^{-5}$, confirming that doubling the loss term does not prevent global collapse.
-NPO and NPO+KL behave differently.
-Even at aggressive LR the shifts remain tightly clustered—most layers move $<\!100$ units on pc1 and almost none on pc2—and green squares consistently fall back a little onto the original line segment.
-RLabel shows an intermediate picture: early layers barely move, while late layers fan out as LR or $N$ grow; nonetheless the rays shorten markedly after relearning, indicating that most distortion is still recoverable.
-Task complexity amplifies divergence.
-On Qwen-2.5-7B the GA rays explode more quickly (Figure 13c,f,i), spanning thousands of pc1 units and driving pc2 to extreme negative values—the hallmark of a high-order, multi-layer perturbation predicted by our theory in Section 4.
-Conversely, NPO’s clusters expand only modestly, even under the same LR, and contract once relearning is applied.
-Taken together, PCA-Shift complements PCA-Similarity: Similarity captures angular mis-alignment, while Shift quantifies translational drift.
-Their joint reading confirms that GA (with or without GD/KL) is prone to large, irreversible representational displacements, whereas NPO variants and, to a lesser extent, RLabel confine shifts to a regime that remains correctable—consistent with the reversible versus irreversible forgetting boundary observed in our utility experiments.
-A.6.2 Centered Kernel Alignment Analysis
-Figures 15– 17 report layer-wise linear CKA between the original model and its unlearned / relearned counterparts.
-Across both Yi-6B and Qwen-2.5-7B, GA again stands out: as LR or $N$ grows its CKA curve close to zero in the final third of the network and never returns, revealing a deep sub-space fracture that matches the irreversible PCA trends.
-GA+GD and GA+KL modestly attenuate this dip, but still fail to restore full alignment after relearning.
-By contrast, NPO and NPO+KL keep CKA higher GA through almost all layers—even under $3{\times}10^{-5}$ or $N{=}100$—and relearning lifts the few layers back to baseline, confirming their perturbations are lower than GA series.
-RLabel occupies an intermediate position: as LR or N increases, its CKA curve drops rapidly—mirroring GA’s behavior—and ultimately exhibits irreversible forgetting.
-Task complexity does not change the ordering but amplifies the gaps: on the math-heavy Qwen benchmark GA’s tail layers fall to almost zero at high LR, whereas NPO keeps higher than GA.
-Taken together with the PCA-Shift results, CKA shows that only GA-style objectives consistently destroy the encoder–decoder sub-space, while NPO families maintain higher stability than GA series and RLabel induces a moderate, recoverable tilt.
-A.6.3 Fisher Information Analysis
-[…]